### v4

# Learning Decision-Focused Uncertainty Sets in Robust Optimization

Abstract
We propose a data-driven technique to automatically learn the uncertainty sets in robust optimization. Our method reshapes the uncertainty sets by minimizing the expected performance across a parametric family of problems, subject to guaranteeing constraint satisfaction. Our approach is very flexible, and can learn a wide variety of uncertainty sets while preserving tractability. We solve the constrained learning problem using a stochastic augmented Lagrangian method that relies on differentiating the solutions of the robust optimization problems with respect to the parameters of the uncertainty set. Due to the nonsmooth and nonconvex nature of the augmented Lagrangian function, we apply the nonsmooth conservative implicit function theorem to establish convergence to a critical point, which is a feasible solution of the constrained problem under mild assumptions. Using empirical process theory, we show finite-sample probabilistic guarantees of constraint satisfaction for the resulting solutions. Numerical experiments show that our method outperforms traditional approaches in robust and distributionally robust optimization in terms of out-of-sample performance and constraint satisfaction guarantees.
1 Introduction
Over the past years, robust optimization (RO) has become a widely adopted efficient tool for decision-making under uncertainty. A robust approach first defines an uncertainty set where the uncertain vector lives, and subsequently hedges its decision against any worst-case realization in this set (Ben-Tal and Nemirovski,, 2008; Ben-Tal et al.,, 2009; Bertsimas et al.,, 2011; Bertsimas and den Hertog,, 2022). To attain proper solution quality and practical robustness the choice of uncertainty set is paramount.
Traditionally, uncertainty sets are shaped and sized to guarantee they contain an appropriately large fraction of the probability mass of the unknown distribution generating the samples (i.e., high coverage). The main appeal of this approach is that the uncertainty set is independent of the subsequent optimization problem considered. In this vein, many approaches to designing uncertainty sets assume structural information of the underlying distributions, and rely on these a priori assumptions to build guarantees of constraint satisfaction (Ben-Tal and Nemirovski,, 2000; Bertsimas et al.,, 2011; Bandi and Bertsimas,, 2012). In particular, a large amount of existing literature studies how to bound the probability of constraint satisfaction based on the size and shape of the uncertainty sets (Li et al.,, 2012; Guzman et al.,, 2017; Bertsimas et al.,, 2021), where the aforementioned structural information is exploited, but the optimization problem itself is not leveraged. By decoupling the design of the uncertainty set from the subsequent optimization problem, the resulting theoretical guarantees can often be quite conservative (Roos and den Hertog,, 2020; Gao and Kleywegt,, 2023).
The trend of constructing uncertainty sets separately from the subsequent optimization problem has continued in the current era of data-driven RO, where the explosion in the availability of data has lead to new paradigms where uncertainty sets are designed directly from data (Bertsimas et al.,, 2018). Whether by combining a priori assumptions with the confidence regions of statistical hypothesis tests on the data (Bertsimas et al.,, 2018), or by approximating a high-probability-region using quantile estimation (Hong et al.,, 2021), significant attention has been given to the probability distribution generating the data and less so to the subsequent optimization problem. More recently in distributionally robust optimization (DRO), the constructed ambiguity set of distributions has historically been required to contain the true data-generating distribution with a certain level of confidence (Esfahani and Kuhn,, 2018). Again, the intuitive appeal of this approach is that the design of the ambiguity set can be carried out in isolation of the subsequent optimization problem.
Unfortunately, decoupling the uncertainty set construction from the subsequent optimization problem in a two-step approach may lead to very conservative decisions. In DRO with Wasserstein ambiguity sets, the two-step approach suffers from the curse of dimensionality as their radius must scale with the dimension of the uncertainty (Weed and Bach,, 2019). On the other hand, one can overcome the curse of dimensionality by exploiting Lipschitz conditions on the objective function of the subsequent optimization problem when designing the ambiguity set (Gao,, 2023). The situation in DRO with $f$-divergence ambiguity sets is even more striking. For instance, Kullback-Leibler ambiguity sets around an empirical distribution can structurally not contain any continuous distribution however large its radius is chosen. Consequently, if the data-generating distribution is continuous a two-step approach is not possible here. Nevertheless, the Kullback-Leibler can be shaped to ensure that the solution has optimal statistical properties when one is willing to exploiting the fact that the subsequent optimization problem has a continuous loss function (Van Parys et al.,, 2021). In both cases, significant reduction in conservatism can be achieved by designing the uncertainty set not only based on the distributions generating the data, but also by exploiting structural properties of the subsequent optimization problem.
The drive for even better performance in RO, then, leads us to a new direction: a method of constructing uncertainty sets that deviates from the conventional decoupled approach and explicitly takes into account the optimization problem as well. This idea has been around for multiple years; as stated by (Ben-Tal et al.,, 2009, page 34), the requirement of containing a high probability mass is only a sufficient (and often restrictive and not needed) condition to guarantee high probability of constraint satisfaction. However, standardized methodology to exploit the notion have seldom been developed. The customary restriction, in essence, allows the uncertainty set developed to hold for all classes of optimization problems, which brings about high levels of conservatism. However, the uncertainty set need only provide high solution quality for the desired optimization problems, and in turn need only include worst-case realizations of uncertainty for these instances.
In this work, we propose a new approach to calibrate uncertainty sets for a particular family of parametric optimization instances. This is a common scenario in many applications, including inventory management, where we solve similar optimization problems with varying initial inventory levels (i.e., the parameters), while satisfying uncertain demand with high probability. Our method minimizes an expected objective over a family of parametric problems, while ensuring data-driven probabilistic guarantees enforced as a conditional value-at-risk constraint (Rockafellar and Uryasev,, 2002). We construct the uncertainty set, which is the family-specific region of worst-case realizations, by learning its shape and size automatically. Specifically, we develop a stochastic augmented Lagrangian approach, relying on nonsmooth implicit differentiation techniques (Agrawal et al., 2019b, ; Agrawal et al., 2019a, ; Bolte et al.,, 2021; Bertrand et al.,, 2022) to compute the derivative of the solution of robust optimization problem with respect to the shape and size of the uncertainty sets. By incorporating the problem objective and the data-driven constraint satisfaction into the learning problem, we control the trade-off between performance and robustness. Our approach offers significant performance improvements over traditional techniques constructing uncertainty sets with high probability mass requirements, and is more flexible than data-driven RO methods that only tune the radius of the uncertainty sets (often via grid search), and not its shape.
1.1 Our contributions
In this work, we present a new approach to automatically learn the uncertainty sets in RO. We obtain high-quality solutions across a family of parametric optimization problems while ensuring probabilistic guarantees of constraint satisfaction. By directly taking into account the quality of the RO solution, the sets found are decision-focused and eliminate the unnecessary conservatism common in traditional data-driven RO methods.
  • We minimize the expectation of the objective function over a family of parametric RO problems, while ensuring an appropriate probability level of constraint satisfaction. The latter is formulated as constraint on the conditional-value-of-risk ($\mathop{\bf CVaR}$) of the problem family’s uncertain constraints.
  • The aforementioned objective and constraints rely on the solutions of the parametric RO problems. The overall formulation is thus a constrained bi-level problem, with the inner level represented by the parametric RO problems.
  • We develop a stochastic augmented Lagrangian algorithm to solve the constrained learning problem to find the optimal parameters of the uncertainty set. Due to the potential nonsmooth, nonconvex, and therefore nondifferentiable nature of the constructed problem, we show path-differentiability of the Lagrangian, and apply nonsmooth implicit differentiation. We show convergence to a critical point under step-size and boundedness conditions, and establish the critical point as a feasible solution of the constrained problem under mild assumptions.
  • Using empirical process theory, we give finite sample probabilistic guarantees of constraint satisfaction for the resulting solution.
  • We benchmark our method on examples in portfolio optimization and multi-stage inventory management, outperforming data-driven robust and distributionally robust approaches in terms of out-of-sample objective value and probabilistic guarantees. The code is available at https://github.com/stellatogrp/lropt_experiments.
1.2 Related work
Data-driven robust optimization.
Data-driven robust optimization has gained wide popularity, by constructing uncertainty sets through data-driven approximations of the unknown data-generating distribution. Using hypothesis testing, Bertsimas et al., (2018) pair a priori assumptions on the distribution with different statistical tests, and obtain various uncertainty sets with different shapes, computational properties, and modeling power. Similar approaches construct data-driven sets using quantile estimation (Hong et al.,, 2021), type-$\infty$ Wasserstein distance (Bertsimas et al.,, 2022), intersections of multiple single-measure-based ambiguity sets (Tanoumand et al.,, 2023), and deep learning clustering techniques (Goerigk and Kurtz,, 2023). All these methods rely on a two-step procedure which separates the construction of the uncertainty set from the resulting robust optimization problem, and often enforces a high coverage requirement on the selected uncertainty sets. As noted by Ben-Tal et al., (2009), however, this requirement may be unnecessary. Recently, Gupta, (2019) proves that while convex problems require DRO ambiguity sets with high coverage to guarantee constraint satisfaction, traditional DRO ambiguity sets for concave problems are significantly larger than the asymptotically optimal sets. In order to find non coverage-based and potentially smaller uncertainty sets that offer improved performance for targeted robust optimization problems, we turn to ideas in differentiable optimization. Recent developments in differentiable convex optimization have enabled a popular end-to-end approach of training a predictive model to minimize loss on a downstream optimization task (Donti et al.,, 2017; Demirović et al.,, 2019; Wilder et al., 2019a, ; Wilder et al., 2019b, ; Cameron et al.,, 2022; Elmachtoub and Grigas,, 2022), for which we also point to the survey by Kotary et al., (2021). Our approach adapts this idea, training instead the shape of the uncertainty set on its performance on parametric robust optimization problems. While developing this work, we became aware of a concurrent work with a similar goal of fitting decision-focused uncertainty sets. In their work, Schuurmans and Patrinos, (2023) construct DRO ambiguity sets which constrain only the worst-case distribution along the direction in which the expected cost of an approximate solution increases most rapidly. Our work, however, considers a broader class of optimization problems and uncertainty sets, and can directly tackle joint chance constraints using hyperparameter tuning.
Contextual optimization.
Contextual optimization and end-to-end approaches, beginning from the Smart Predict, then Optimize framework (Elmachtoub and Grigas,, 2022) have been shown to improve performance as opposed to traditional two-stage approaches, where the model is trained separately from the optimization problem (Costa and Iyengar,, 2023).  Costa and Iyengar, (2023) propose an end-to-end distributionally robust approach for portfolio construction, where they integrate a prediction layer that predicts asset returns with data on financial features and historical returns, and use these predictions in a decision layer that solves the robust optimization problem. Concurrent to our work, Chenreddy and Delage, (2024) propose an end-to-end training algorithm to produce contextual uncertainty sets that achieve a certain conditional coverage. This is to decrease the conservatism of their previous two-stage approach, conditional robust optimization, where they learn a contextually adapted uncertainty set known to contain with high probability the realized uncertainty (Chenreddy et al.,, 2022). We, however, adopt a different notion for context parameters, denoting them family parameters that parameterize different instances of the downstream optimization problem, instead of the uncertain parameter. Our learned uncertainty set therefore does not learn a direct mapping from the context parameters to the uncertainty set (as in contextual and conditional RO (Sadana et al.,, 2023)), as it may be computationally expensive and hard to interpret. We do not enforce coverage constraints, which may introduce unnecessary conservatism, and instead freely shape the uncertainty sets based on their performance on the entire parametralized family of robust problems. As pointed out in the recent survey by Sadana et al., (2023), current work in the end-to-end framework seldom focus on uncertainty in the downstream optimization problem, especially in the constraints. Our approach thus fills this gap, where the process of constructing the uncertainty set is linked in an end-to-end fashion to the solution of the inner robust optimization problem. Our formulation is flexible, as it allows for a wide variety of uncertain constraints and uncertainty sets.
Modeling languages for robust optimization.
Several open-source packages facilitate building models for RO, including as ROME (Goh and Sim,, 2009) in Matlab, RSOME (Xiong and Chen,, 2021) in Python, JuMPeR (Dunning et al.,, 2017) in Julia, and ROmodel (Wiebe and Misener,, 2022) for the Pyomo modeling language in Python. These packages mostly support RO problems with affine uncertain constraints, which are solved via reformulations and/or cutting plane procedures. Schiele et al., (2023) recently introduced a Python extension of cvxpy (Diamond and Boyd,, 2016) for solving the larger class of convex-concave saddle point problems, by automating a conic dualization procedure reformulating min-max type problems into min-min problems which can then be solved efficiently (Juditsky and Nemirovski,, 2022). In the paper, they describe a set of rules termed disciplined saddle programming (DSP) detailing when a saddle point optimization problem can be equivalently formulated as convex optimization ones, and efficiently be solved. In this work, we build a Python extension to cvxpy tailored to formulating and solving RO problems. While some RO problems can be described in terms of the DSP framework, our library augments DSP capabilities in RO by supporting max-of-concave uncertain constraint and structure to intuitively describe well known uncertainty sets for ease of use. Finally, our method automatically learns the uncertainty set through differentiable optimization via a tight integration with the works by Agrawal et al., 2019a ; Agrawal et al., 2019b .
Differentiable convex optimization.
There has been extensive work on embedding differentiable optimization problems as layers within deep learning architectures (Amos and Kolter,, 2017). Agrawal et al., 2019a propose an approach to differentiating through disciplined convex programs, and implement their methodology in cvxpy. Specifically, they implement differentiable layers for disciplined convex programs in PyTorch and TensorFlow 2.0, in the package cvxpylayers. They convert the convex programs into conic programs, then implicitly differentiate the residual map of the homogeneous self-dual embedding associated with the conic program. These developments have fueled the recent advances in end-to-end approaches, and is also the tool of choice in this work.
Nonsmooth implicit differentiation.
Traditional techniques in differentiable convex optimization rely on the implicit function theorem, which requires derivatives to exist almost everywhere. Recently, a nonsmooth variant has been established, using conservative Jacobians, provided that a nonsmooth form of the classical invertibility condition is fulfilled (Bolte et al.,, 2021). As a key ingredient, definable and locally Lipschitz continuous functions are shown to admit conservative Jacobians, which follow the chain rule, and are thus compatible with backpropagation techniques in stochastic gradient descent (Davis et al.,, 2018; Bolte et al.,, 2021). These techniques can be used to obtain convergence for nonsmooth problems.
Automatic hyperparameter selection.
Implicit differentiation techniques can greatly accelerate tuning the regularization parameters in regression and classification problems (Bertrand et al.,, 2020, 2022; Blanchet et al.,, 2019). However, most literature focuses on tuning hyperparameters of the objective function in regression and classification tasks, such as hyper-parameters for Lasso regularization (Bertrand et al.,, 2022). In this work, we tune various hyperparameters of the uncertainty sets in RO using implicit differentiation techniques.
Risk measures.
Value at risk ($\mathbf{VaR}$) and conditional value at risk ($\mathop{\bf CVaR}$) are common risk measures for quantifying costs encountered in the tails of distributions. A $\mathbf{VaR}_{\eta}$ constraint is a chance constraint that measures the maximum cost incurred with respect to a given confidence level $\eta$, but while intuitive, the function is often intractable, and does not quantify costs beyond the found threshold value. Conditional value at risk, on the other hand, measures the expected cost given that the cost exceeds $\mathbf{VaR}$, and is known to have a tractable formulation (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002). Since its development, $\mathop{\bf CVaR}$ optimization has been a common method for controlling risk in portfolio optimization and option hedging problems, and has also been explored in risk-constrained Markov decision problems, where the risk is presented as a constraint on the $\mathop{\bf CVaR}$ of the cumulative cost (Chow et al.,, 2015). The use of $\mathop{\bf CVaR}$ constraints to substitute chance constraints has also been explored in reliability engineering, where the $\mathop{\bf CVaR}$ constraint has been given a different name: buffered failure probability (Rockafellar and Royset,, 2010). A related concept for controlling risk is the buffered probability of exceedance (bPOE) (Mafusalov and Uryasev,, 2018), which is equivalent to 1 minus the inverse of $\mathop{\bf CVaR}$, and has a level set equivalence with $\mathop{\bf CVaR}$. In our work, we constrain risk, which in RO is the probability of constraint violation, by constraining the $\mathop{\bf CVaR}$ of the uncertain constraint functions.
1.3 Layout of the paper
In Section 2, we introduce the notion of the reshaping parameters with a motivating example. In Section 3, we describe the data-driven problem and the algorithm for learning reshaping parameters, and prove its convergence. In Section 4, we discuss the probabilistic guarantees implied by the aforementioned formulations. In Section 5, we present various numerical examples problems. For completeness, we give example convex reformulations of robust optimization problems for common uncertainty sets, with our reshaped parameters, in the appendices.
2 The robust optimization problem
We consider a family of parametric robust optimization (RO) problems, indexed by the family parameter $y\in{\mbox{\bf R}}^{p}$, where each instance takes the form:
$x(\theta,y)\in\begin{array}[t]{ll}\mathop{\rm argmin}&f(x)\\ \mbox{subject to}&g(x,u,y)\leq 0\quad\forall u\in{\mathcal{U}}(\theta),\end{array}$ (1)
where $x\in{\mbox{\bf R}}^{n}$ is the optimization variable, $u\in{\mbox{\bf R}}^{m}$ is the uncertain parameter, $f:{\mbox{\bf R}}^{n}\to{\mbox{\bf R}}$ is the convex objective function, and $g:{\mbox{\bf R}}^{n}\times{\mbox{\bf R}}^{m}\times{\mbox{\bf R}}^{p}\to{\mbox{\bf R}}$ is the uncertain constraint which we assume to be the maximum of functions concave in $u$ and convex in $x$ and $y$e.g., $g(x,u,y)=\max_{l=1,\dots,L}g_{l}(x,u,y)$. This structure allows us to express $L$ joint uncertain constraints using the maximum of each component $l$. Problem (1) guarantees that, for a given $y$, the solution $x(\theta,y)$ satisfies the constraints for any value of the uncertain parameter $u$ in ${\mathcal{U}}(\theta)\subseteq{\mbox{\bf R}}^{m}$, a convex uncertainty set that depends on the set parameter $\theta\in{\mbox{\bf R}}^{q}$. For example, $\theta$ can represent the center, size, and shape of ${\mathcal{U}}(\theta)$. In this setting, the family parameter $y$ represents information available when we make our decisions (e.g., current allocations in portfolio optimization). The uncertain parameter $u$, however, is not known (e.g., future returns in portfolio optimization). Importantly, we do not assume that $y$ affects $u$, which is common in several practical applications (e.g., when our portfolio allocations do no affect future market returns).
Here, the set ${\mathcal{U}}(\theta)$ does not depend on individual $y$ as this would require either learning a different uncertainty set for every instance of the family, which may be costly, or learning a mapping from $y$ to the uncertainty set (as in conditional RO (Sadana et al.,, 2023)), which may be hard to interpret. Instead, we construct a fixed set that works for all realizations of $y$ encountered in target application.
Our goal is to find a set parameter $\theta$ in a decision-focused manner, i.e., such that the solutions $x(\theta,y)$ perform well in terms of objective value and constraint satisfaction guarantees across the family of problems. Specifically, we choose the uncertainty set ${\mathcal{U}}(\theta)$ such that the solution $x(\theta,y)$ from (1) implies that the uncertain constraint is satisfied with probability $1-\eta$ with $\eta>0$, across all instances of problem (1):
${\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq 0)\geq 1-\eta.$ (2)
Here, ${\mathbf{P}}_{y}$ is the distribution of $y$ with domain $\mathcal{D}_{y}\subseteq{\mbox{\bf R}}^{p}$, ${\mathbf{P}}_{u}$ is the distribution of $u$ with domain $\mathcal{D}_{u}\subseteq{\mbox{\bf R}}^{m}$, and ${\mathbf{P}}_{(u,y)}$ is their joint distribution, all of which are unknown. We note that this is an aggregate probabilistic guarantee for the parametric family of problems, and not conditional guarantee which holds for all instances $y$. Whereas such conditional guarantee (Chenreddy and Delage,, 2024) are perhaps more desirable they require stringent assumptions on the distribution ${\mathbf{P}}_{y}$ which are typically hard to verify in practice. We can formulate constraint (2) in terms of the corresponding value at risk, i.e.,
$\mathop{\bf VaR}(g(x(\theta,y),u,y),\eta)=\inf\{\gamma\mid{\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq\gamma)\geq 1-\eta\}\leq 0.$
Unfortunately, the value at risk function often leads to intractable formulations (Uryasev and Rockafellar,, 2001). To get a tractable approximation, we can adopt the conditional value at risk (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002), defined as
$\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)=\inf_{\alpha}\left\{\frac{{\mbox{\bf E}}_{(u,y)}(g(x(\theta,y),u,y)-\alpha)_{+}}{\eta}+\alpha\right\},$ (3)
where $(a)_{+}=\max\{a,0\}$. It is well known (Uryasev and Rockafellar,, 2001) that, for any $x\in{\mbox{\bf R}}^{n}$, the relationship between these probabilistic guarantees of constraint satisfaction is
$\mathop{\bf CVaR}(g(x,u,y),\eta)\leq 0~{}~{}\Longrightarrow~{}~{}\mathop{\bf VaR}(g(x,u,y),\eta)\leq 0~{}~{}\Longleftrightarrow~{}~{}{\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta.$ (4)
Therefore, if the solution $x(\theta,y)$ of (1) satisfies $\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)\leq 0$, we have the desired probabilistic guarantee (2). This constraint, together with the goal of minimizing the expected objective value, are constructed into an evaluation metric for updating $\theta$; see Figure 1. Typically, by constructing uncertainty sets that contain at least $1-\eta$ probability mass, i.e.${\mathbf{P}}_{u}(u\in{\mathcal{U}}(\theta))\geq 1-\eta$, we can ensure that any feasible solution of (1) implies a probabilistic guarantee (2). However, this sufficient condition may lead to overly conservative solutions (Ben-Tal et al.,, 2009, page 34), as it works for any RO problem and any constraint using the uncertainty set ${\mathcal{U}}(\theta)$. In this work, instead, we avoid such conservatism by taking into account the cost of the robust solutions $x(\theta,y)$ while constructing the uncertainty sets.
Figure 1: Our high-level procedure. We evaluate the performance of the decision $x(\theta,y)$ in terms of the expected objective value and constraint satisfaction. The evalution metric is defined in Section 3.2.
2.1 Motivating example
We consider a 2-product newsvendor problem where each day, the vendor orders $x\in{\mbox{\bf R}}^{2}$ products at price $k\in{\mbox{\bf R}}^{2}$, and we allow fractional values for simplicity. The products are sold at price $p\in{\mbox{\bf R}}^{2}$, until either the demand $u\in{\mbox{\bf R}}^{2}$ or inventory $x$ is exhausted. The objective function to minimize is the sum of the ordering cost minus the revenue, $f(x,u,k,p)=k^{T}x-p^{T}\min\{x,u\}=k^{T}x+\max(-p_{1}x_{1}-p_{2}x_{2},-p_{1}x_{1}-p_{2}u_{2},-p_{1}u_{1}-p_{2}x_{2},-p_{1}u_{1}-p_{2}u_{2}).$ We obtain the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t\\ \mbox{subject to}&f(x,u,k,p)\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &x\geq 0.\end{array}$ (5)
We parametrize the problem (5) by $y=(k,p)$, the buying and selling prices of the products, and would like to construct uncertainty sets without knowing the distributions of $y$ and $u$.
Problem setup.
The uncertain parameter $u$ is distributed as a log-normal distribution, with parameters
$\mu=\left[\begin{array}[]{l}1.1\\ 1.7\end{array}\right],\quad\Sigma=\left[\begin{array}[]{ll}\phantom{-}0.6&-0.3\\ -0.3&\phantom{-}0.1\end{array}\right],$
and we have access to $50$ independent samples. We generate 10 values of $(k,p)$, where each $k=[4,5]+\delta_{1}$ and $p=k+\max\{0,\delta_{2}\}$, with $\delta_{1},\delta_{2}$ standard normal random variables. We repeat the data generation for $u$ and $y$ detailed above, for 15 repetitions of the experiment. Below, we show the uncertainty sets obtained for one such repetition.
Figure 2: Top (bottom) row: mean-variance (reshaped) uncertainty set, depicted in red, corresponding to the empirical probability of constraint violation $\hat{\eta}$. White dots: training data points. Green dot: training data mean. Middle row: out-of-sample trade-off curves for the objective value vs. empirical probability of constraint violation, averaged across all problem instances and all repetitions of the experiment. The shaded regions represent the 0.3 to 0.7 quantiles of the repetitions. Each marker corresponds to an $\rho$ value. The uncertainty sets depicted above and below achieves the probabilities of constraint violation $\hat{\eta}$ closest to the red dotted lines, for one repetition of the experiment.
Mean-Variance uncertainty set.
Standard methods in RO methods construct uncertainty sets based on the empirical mean $\hat{\mu}$ and covariance $\hat{\Sigma}$ of the uncertainty,
${\mathcal{U}}{(\theta)}=\{u=\hat{\mu}+\hat{\Sigma}^{1/2}z\mid\|z\|_{2}\leq\rho\}=\{b^{\rm mv}+A^{\rm mv}z\mid\|z\|_{2}\leq\rho\},$ (6)
where $\theta=(A^{\rm mv},b^{\rm mv},\rho)$. The parameter $\rho$ represents the size of the uncertainty set, and, for one set of data, we have
$A^{\rm mv}=\hat{\Sigma}^{1/2}=\left[\begin{array}[]{ll}\phantom{-}3.08&-0.83\\ -0.83&\phantom{-}1.49\end{array}\right],\quad b^{\rm mv}=\hat{\mu}=\left[\begin{array}[]{l}3.96\\ 5.82\end{array}\right].$
Decision-focused (reshaped) uncertainty set.
Consider now the same set as in (6) with different shape, i.e., ${\mathcal{U}}(\theta)=\{u=b^{\rm re}+A^{\rm re}z\mid\|z\|_{2}\leq\rho\}$, where $\theta=(A^{\rm re},b^{\rm re},\rho)$ and
$A^{\rm re}=\left[\begin{array}[]{ll}0.12&2.05\\ 2.09&0.62\end{array}\right],\quad b^{\rm re}=\left[\begin{array}[]{l}2.7\\ 4.8\end{array}\right].$
Comparison.
In Figure 2, we compare the two uncertainty sets where we vary $\rho$ while $A^{\rm st},b^{\rm st}$ and $A^{\rm re},b^{\rm re}$ are fixed. For each value of $\rho$ we compare the average out-of-sample objective value at the optimizers $x(\theta,y)$ against the out-of-sample empirical probability of constraint violation. The average is taken across problem instances $y$’s, testing datasets of $N=450$ uncertainty realizations, and over all repetitions of the experiment. We note that the reshaped uncertainty set achieves data-driven solutions that give, on average, better trade-offs. While the mean-variance uncertainty set conforms to the shape of data, it might be conservative because it does not take into account the structure of the optimization problem. In contrast, the reshaped uncertainty set gives better (lower) out-of-sample costs for the same probability of constraint satisfaction, even though it does not conform as well to the shape of the data, and covers a lower percentage of the data points.
In the next sections, we introduce an automatic technique to find such sets with better performance, by tuning their shape and size such that the problem objective is minimized while guaranteeing constraint satisfaction across the family of parametric RO problems (1).
3 Learning the uncertainty set
3.1 Constrained learning formulation
Consider a RO problem (1). We have a dataset of $N$ independent samples of the uncertain parameter $u$, $U^{N}=\{u^{i}\}_{i=1}^{N}\subseteq\mathcal{D}_{u}$, governed by the product distribution ${\mathbf{P}}^{N}_{u}$, and a dataset of $J$ independent samples of the family parameter $y$, $Y^{J}=\{y^{j}\}_{j=1}^{J}\subseteq\mathcal{D}_{y}$, governed by the product distribution ${\mathbf{P}}_{y}^{J}$. We define the combined samples $(u^{i},y^{j})$ for $i=1,\dots,N$ and $j=1,\dots,J$, and the combined dataset $U^{N}\times Y^{J},$ with product distribution defined as ${\mathbf{P}}^{N\times J}$. Our main requirement is to identify a set parameter $\theta$ for which the solutions $x(\theta,y)$ to problem (1) with uncertainty set $\mathcal{U}(\theta)$, satisfy the following finite-sample probabilistic guarantee
${\mathbf{P}}^{N\times J}\left\{{\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq 0)\geq 1-\eta\right\}\geq 1-\beta,$ (7)
where $\beta>0$ is the target confidence level across draws of the dataset $U^{N}\times Y^{J}$. Instead of working with the probability ${\mathbf{P}}_{(u,y)}$, which often leads to intractable formulations (Uryasev and Rockafellar,, 2001) as discussed in Section 2, we adopt a tractable approximation based on the conditional value at risk (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002) defined in (3), obtaining the constraint
$\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)\leq\kappa\quad\iff\quad\inf_{\alpha}\left\{\frac{{\mbox{\bf E}}_{(u,y)}(g(x(\theta,y),u,y)-\alpha)_{+}}{\eta}+\alpha\right\}\leq\kappa,$
where $\kappa>0$ is a threshold value that we define in Section 4 to obtain the finite sample guarantees (7). Our goal is to find the set parameter $\theta$ that minimizes the average objective of the RO problem across realizations of the family parameter $y$, which can be written as
$\begin{array}[]{ll}\mbox{minimize}&{\mbox{\bf E}}_{y}\left(f(x(\theta,y))\right)\\ \mbox{subject to}&{\mbox{\bf E}}_{(u,y)}\left((g(x(\theta,y),u,y)-\alpha)_{+}/\eta+\alpha\right)\leq\kappa,\\ \end{array}$ (8)
where $x(\theta,y)$ are solutions of the inner RO problems, defined as in (1). We given in Appendix A examples of parameterized uncertainty sets and the corresponding convex reformulations of (1), which we use to solve for $x(\theta,y)$. The expectation constraint corresponds to the $\mathop{\bf CVaR}$ constraint where $\alpha$ is shifted to become a minimization variable. When the functions $f$ and $g$ are convex in $x$, this shift in $\alpha$ has no impact on the optimal solution (Sarykalin et al.,, 2008). With this formulation, we tune $\alpha$ and the set parameter $\theta$ to improve the quality of the solutions $x(\theta,y)$ of the inner RO problem (1). Solving the training problem (8) is challenging because of the stochasticity in the objective and the constraints, and the nonconvex bi-level structure introduced by the inner RO problem. We address these challenges using a stochastic augmented Lagrangian algorithm.
3.2 Training the uncertainty set using a Stochastic augmented Lagrangian algorithm
As the distributions ${\mathbf{P}}_{u}$ and ${\mathbf{P}}_{y}$ are unknown, we are unable to obtain the true expectation in (8). As such, we work with sampled versions of the expectations. We define a condition on the sampled $\mathop{\bf CVaR}$,
$\widehat{\mathop{\bf CVaR}}(\theta)=\inf_{\alpha}\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\left(\frac{(g(x(\theta,y^{j}),u^{i},y^{j})-\alpha)_{+}}{\eta}+\alpha\right)\leq\kappa,$ (9)
and require this condition to hold for feasible $\theta$. To find such a $\theta$, we introduce finite-sample approximations of the objective and constraint functions of problem (8),
$F(\theta)=\frac{1}{J}\sum_{j=1}^{J}f(x(\theta,y^{j}))\quad\text{and}\quad H(\alpha,\theta)=\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}(g(x(\theta,y^{j}),u^{i},y^{j})-\alpha)_{+}/\eta+\alpha,$
evaluated over the dataset $U^{N}\times Y^{J}$. We use these approximations to define the sampled version of problem (8),
$\begin{array}[]{ll}\mbox{minimize}&F(\theta)\\ \mbox{subject to}&H(\alpha,\theta)+s=\kappa\\ &s\geq 0,\\ \end{array}$ (10)
where $s$ is the non-negative slack variable in ${\mbox{\bf R}}^{d}$ to represent the inequality constraint as an equality. We can now write the augmented Lagrangian function as
$L(\alpha,s,\theta,\lambda,\mu)=F(\theta)$ $+\lambda(H(\alpha,\theta)+s-\kappa)+\frac{\mu}{2}\|H(\alpha,\theta)+s-\kappa\|^{2},$ (11)
where $\lambda$ is the Lagrange multiplier in R, and $\mu$ is the penalty parameter in R. This function $L$ is evaluated over the entire dataset $U^{N}\times Y^{J}$. We denote the estimated augmented Lagrangian over a subset (batch) of data $B=(U\times Y)\subseteq(U^{N}\times Y^{J})$ as $\hat{L}$. Using the stochastic augmented Lagrangian approach outlined in Algorithms 1 and 2, and under mild assumptions, we can find a feasible $\theta$ satisfying condition (9). We also define a slightly weaker optimality condition: an $\epsilon$-stationary point for problem (10). Given $\epsilon>0$, a point $(\alpha,s,\theta)\in{\mbox{\bf R}}^{q+2}$ is called an $\epsilon$-stationary point of (10) if there is a scalar $\lambda\in{\mbox{\bf R}}$ such that
$\|H(\alpha,\theta)+s-\kappa\|_{2}\leq\epsilon,\quad\mathop{\bf dist{}}(0,J_{F}(\theta)+\lambda J_{H+s-\kappa}(\alpha,s,\theta)))\leq\epsilon,$ (12)
where $J_{F}(\theta)$ is the conservative Jacobian (Definition 3.1 below) of the objective function $F$ in the variable $\theta$, $J_{H+s-\kappa}(\alpha,s,\theta)$ denotes the conservative Jacobian of the constraint function in variables $(\alpha,s,\theta)$, and the distance between a vector $x$ and set $\mathcal{X}$ is defined as $\mathop{\bf dist{}}{(x,\mathcal{X})}=\min_{y\in\mathcal{X}}\|x-y\|_{2}$. In addition, given $\lambda$ and $\mu$, a point $(\alpha,s,\theta)\in{\mbox{\bf R}}^{q+2}$ is called a stationary point of the augmented Lagrangian function (11), if $0\in J_{L}(\alpha,s,\theta)$. While useful in our convergence analysis, this stationarity condition is hard to evaluate. That’s why in practical applications, instead of using it to terminate Algorithm 2, we set the number of iterations $t_{\text{max}}$ to a reasonably large value.
Algorithm 1 Safeguarded stochastic augmented Lagrangian algorithm to solve (10)
1:given $\theta^{0}=(A^{\text{init}},b^{\text{init}})$, $\alpha^{0}$, $s^{0}$, $\lambda^{0},\mu^{0},\epsilon\geq 0,\kappa$, $\sigma>1$, $\tau<1$, $H_{\text{best}}=\infty$
2:for $k=1,\dots,k_{\rm max}$ do
3:     $(\alpha^{k},s^{k},\theta^{k})\leftarrow$ minimize $L(\alpha,s,\theta,\lambda^{k-1},\mu^{k-1})$ over $(\alpha,s,\theta)$ via Algorithm 2
4:     if $\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}\leq\epsilon$ then
5:         return $\theta^{k}$triangleright $\epsilon$-stationary point found
6:     else if $\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}\leq\tau\|H_{\text{best}}\|_{2}$ then
7:         $\lambda^{k}\leftarrow\Pi_{[{\lambda^{\rm{min}}},\lambda^{\rm{max}}]}(\lambda^{k-1}+\mu^{k-1}(H(\alpha^{k},\theta^{k})+s^{k}-\kappa)$) $\triangleright$ update Lagrange multipliers
8:         $H_{\text{best}}\leftarrow\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}$
9:         $\mu^{k}\leftarrow\mu^{k-1}$
10:     else
11:         $\mu^{k}\leftarrow\sigma\mu^{k-1}$triangleright update penalty
12:         $\lambda^{k}\leftarrow\lambda^{k-1}$
13:return $\theta^{k_{\rm max}}$
Algorithm 2 Stochastic gradient descent for the inner loop
1:given $\theta^{0}$, $\alpha^{0}$, $s^{0}$, $\lambda,\mu,\{\delta^{t}\}_{t\in\mathbf{N}},U^{N}\times Y^{J}$
2:for $t=1,\dots,t_{\rm max}$ do
3:     sample data $(U^{t}\times Y^{t})\subseteq(U^{N}\times Y^{J})$
4:     $x^{t}(\theta^{t-1},y)\leftarrow$ solve inner RO problem (1) for all $y\in Y^{t}$
5:     $J_{\hat{L}}\leftarrow J_{\hat{L}}(\alpha^{t-1},\theta^{t-1},s^{t-1})$triangleright compute Jacobian of $\hat{L}$ over $U^{t}\times Y^{t}$
6:     $\alpha^{t}\leftarrow\alpha^{t-1}-\delta^{t}v_{\alpha}^{t},\quad\text{with}\quad v_{\alpha}^{t}\in J_{\hat{L}}(\alpha)$
7:     $s^{t}\leftarrow(s^{t-1}-\delta^{t}v_{s}^{t})_{+},\quad\text{with}\quad v_{s}^{t}\in J_{\hat{L}}(s)$
8:     $\theta^{t}\leftarrow\theta^{t-1}-\delta^{t}v_{\theta}^{t},\quad\text{with}\quad v_{\theta}^{t}\in J_{\hat{L}}(\theta)$
9:return $(\alpha^{t},s^{t},\theta^{t})$
3.2.1 Path differentiability and conservative Jacobians
In order to apply Algorithms 1 and 2, we need to establish the existence of $J_{L}$ and $J_{\hat{L}}$, the conservative Jacobians of the augmented Lagrangian functions $L$ and $\hat{L}$. Conservative Jacobians, introduced in (Bolte and Pauwels,, 2019), are a generalized form of Jacobians for functions that are almost everywhere differentiable.
Definition 3.1 (Conservative Jacobian (Bolte et al.,, 2021)).
Given a locally Lipschitz continuous function $F:{\mbox{\bf R}}^{m}\rightarrow{\mbox{\bf R}}^{m}$, we call $J_{F}:{\mbox{\bf R}}^{n}\rightrightarrows{\mbox{\bf R}}^{n\times m}$ its conservative Jacobian if $J_{F}$ has a closed graph, is locally bounded, and is nonempty with
[…]
for any absolutely continuous curve $\gamma$ in ${\mbox{\bf R}}^{n}$. A locally Lipschitz function $F$ that admits a conservative Jacobian $J_{F}$ is called path differentiable.
To obtain the conservative Jacobians $J_{L}$ and $J_{\hat{L}}$, we note the following assumptions.
Assumption 3.1 (Definability).
All functions $f$ and $g_{l}$ in the robust optimization problem (1) are definable (van den Dries and Miller,, 1999; Coste,, 2000; Bolte et al.,, 2021).
The class of definable functions is very wide and includes semi-algebraic functions; for an overview, see the works by Bolte and Pauwels, (2019) and Bolte et al., (2021). For a foundational work, see van den Dries and Miller, (1999). Assuming the definability of the objective and constraint functions, it follows that $F$ and $H$, and subsequently $L$, as compositions of definable pieces, are also definable (Bolte et al.,, 2021).
Assumption 3.2 (Local Lipschitzness).
The Lagrangian function $L(\alpha,s,\theta,\lambda,\mu)$ is locally Lipschitz in variables $\alpha,s,\theta,\lambda,\mu$. A function $F:{\mbox{\bf R}}^{n}\rightarrow{\mbox{\bf R}}^{m}$ is locally Lipschitz if, for each $x\in{\mbox{\bf R}}^{n}$, there exists a neighborhood $\mathcal{V}$ of $x$ such that $F$ is Lipschitz on $\mathcal{V}$.
We also require the augmented Lagrangian to be locally Lipschitz continuous, as definable locally Lipschitz functions are shown to be path differentiable, meaning they admit conservative Jacobians and follow the chain rule (Bolte et al.,, 2021). Admittance of the chain rule is imperative, as $L$ depends on $\theta$ indirectly through the solution $x(\theta,y)$ of the inner robust problem (1). In order to establish the conservative Jacobians of $L$ in the variable $\theta$, we thus need to show the path differentiability of the inner problem and compute $J_{x}(\theta)$, for which we require a uniqueness assumption.
Assumption 3.3 (Uniqueness).
The robust optimization problem (1) has a convex conic reformulation with a unique solution.
We can then make use of the results of Agrawal et al., 2019b and Bolte et al., (2021), which enables us to find conservative Jacobians of optimal solutions for conic problems by differentiating through the KKT optimality conditions. We state the theorem below, with the proof delayed to Appendix C.
Theorem 3.1 (Path differentiability of the RO solution).
Suppose Assumption 3.3 holds. Let $x(\theta,y)$ be the unique solution, for a given input pair $(\theta,y)$, to the convex conic reformulation of the inner optimization problem (1). This solution is path differentiable with nonempty conservative Jacobian $J_{x}(\theta)\neq\emptyset$.
We can then evaluate the conservative Jacobians of the augmented Lagrangian function.
Proposition 3.1 (Conservative Jacobians of the augmented Lagrangian (11)).
Suppose Assumptions 3.1 and 3.2 hold, and for each given input pair $(\theta,y)$, the solution $x(\theta,y)$ of the optimization problem (1) is path differentiable with conservative Jacobian $J_{x}(\theta)$. Then, the Lagrangian function $L(\alpha,s,\theta,\lambda,\mu)$, from equation (11), and its batched version $\hat{L}(\alpha,s,\theta,\lambda,\mu)$, are path differentiable with conservative Jacobians $J_{L}(\alpha,s,\theta)$ and $J_{\hat{L}}(\alpha,s,\theta)$.
The specific forms of the conservative Jacobians established in Proposition 3.1 are given in Appendix B.
3.2.2 Convergence analysis
To show convergence of the augmented Lagrangian algorithm, given in Algorithm 1, we require the convergence of the inner loop, Algorithm 2. The convergence of Algorithm 2 relies on the path-differentiability of the Lagrangian function, as well as the boundedness of the iterates and the Lagrangian multiplier $\lambda$.
Assumption 3.4 (Boundedness).
There exists a constant $\alpha^{\rm max}>0$ and compact set $\mathcal{D}_{\alpha}\subset{\mbox{\bf R}}$, such that, for $\alpha^{0}\in\mathcal{D}_{\alpha}$, $\|\alpha^{t}\|\leq\alpha^{\rm max}$ almost surely. Similarly, there exist constants $\theta^{\rm max},s^{\rm max},x^{\rm max},y^{\rm max}>0$, compact sets $\mathcal{D}_{\theta}\subset{\mbox{\bf R}}^{q}$, $\mathcal{D}_{s}\subset{\mbox{\bf R}}$, and $\mathcal{X}\subset{\mbox{\bf R}}^{n}$, such that, for $\theta^{0}\in\mathcal{D}_{\theta}$, $s^{0}\in\mathcal{D}_{s}$, and $y\in\mathcal{D}_{y}$, we have $\|\theta^{t}\|\leq\theta^{\rm max}$, $\|s^{t}\|\leq s^{\rm max}$, and $x(\theta^{t},y)\in\mathcal{X}=\{x\in{\mbox{\bf R}}^{n}\mid\|x\|\leq x^{\rm max}\}$ almost surely. In addition, the optimal Lagrangian multiplier $\lambda^{\star}\in[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$.
Lastly, we require that the step sizes are nonnegative, square summable, but not summable.
Assumption 3.5 (Step sizes).
The step sizes $\{\delta^{t}\}_{t\in\mathbf{N}}$ are chosen such that
$\sum_{t=1}^{\infty}\delta^{t}=\infty,\quad\sum_{t=1}^{\infty}(\delta^{t})^{2}<\infty,\quad\delta^{t}=o(1/\log(t)).$
Under these assumptions, by Theorem 6.2 of Davis et al., (2018) and Theorem 3 of Bolte et al., (2021), the sequence $\{\alpha^{t},s^{t},\theta^{t}\}$ in Algorithm 2 converges to a stationary point of the Lagrangian function (11) with arbitrary $\lambda$ and $\mu$.
We note that $\lambda^{k}$ and $\mu^{k}$ in Algorithm 1 are potentially unbounded, in which case they do not converge. The stationary point $(\alpha^{k},s^{k},\theta^{k})$ of the augmented Lagrangian function for any $\lambda^{k}$ and $\mu^{k}$ is, therefore, not guaranteed to be a feasible solution of the $\mathop{\bf CVaR}$-constrained problem (10). However, with the use of the safeguards $[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$, we can guarantee that the limit point of the augmented Lagrangian algorithm is at least a stationary point of constraint violation (Birgin and Martinez,, 2009, Theorem 6.3). More specifically, below we show that the algorithm finds a feasible point of condition (9) under mild assumptions, and otherwise converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$. The proof is delayed to Appendix D.
Theorem 3.2.
Suppose Assumptions 3.1 to 3.5 hold. If the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, then the sequence of updates $\{\theta^{k}\}$ in Algorithm 1 converges almost surely to a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition (9) as $k\to\infty$. Otherwise, $\{\alpha^{k},s^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$.
Note that a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition (9) also implies a feasible point of the contrained problem (10). In practice, with a finite number of iterations and some $\epsilon>0$, we obtain an $\epsilon$-stationary solution of (10) satisfying condition (12).
4 Probabilistic guarantees of constraint satisfaction
In the previous section, we have proved convergence to a feasible solution of the sampled problem (10), which does not exactly correspond to the probabilistic guarantees of constraint satisfaction (7). To arrive at our desired guarantees, where the inner probability is with respect to the unknown joint distribution ${\mathbf{P}}_{(u,y)}$, we need to analyze the generalization behavior of the robust solutions on out-of-sample data. We begin with a boundedness assumption on the function inside the expectation from the $\mathop{\bf CVaR}$ definition, which appears in (8).
Assumption 4.1.
The function $g(x(\theta,y),u,y)-\alpha)_{+}/\eta+\alpha$ maps to values in $(-c^{\rm min},c^{\rm max})$, where $0\leq c^{\rm min},c^{\rm max}<\infty$, for all $\theta,\alpha,u,y$ in its support.
Let $C=c^{\rm min}+c^{\rm max}$ and define the normalized function
$\psi_{z}(u,y)=\frac{1}{C}\left(c^{\rm min}+\frac{(g(x,u,y)-\alpha)_{+}}{\eta}+\alpha\right),$ (13)
where $z=(x,\alpha)\in\mathcal{Z}\subset{\mbox{\bf R}}^{n+1}$. Note that when Assumption 3.4 holds, we have $\mathcal{Z}=\mathcal{X}\times\mathcal{D}_{\alpha}$ being a compact set. The function $\psi_{z}$ then maps to values in $(0,1)$, and we consider the entire function class
$\Psi=\{\psi_{z}\}_{z\in\mathcal{Z}},$
which we use to acquire the finite sample probabilistic guarantee in Theorem 4.1.
Theorem 4.1 (Finite sample probabilistic guarantee).
[…]
$\sup_{Q}N(\delta,\Psi,L_{2}(Q))\leq\left(\frac{K}{\delta}\right)^{V},$ (14)
for every $0<\delta<K$. Then, when Algorithm 1 converges to an $\epsilon$-stationary point of (10), we have the finite sample probabilistic guarantee
${\mathbf{P}}^{N\times J}\left({\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta\quad\forall x\in\mathcal{X}\right)\geq 1-\beta,$ (15)
where $\beta=(D\sqrt{N}C\tau/\sqrt{V})^{V}\exp(-2NC^{2}\tau^{2})$, $D$ is a constant that depends only on $K$, and $\tau\leq(c^{\min}+\kappa+\epsilon)/C$.
Proof.
By the definition of the function class $\Psi$ and Assumptions 3.4 and 4.1, we can apply Theorem 2.14.9 by van der Vaart and Wellner, (1996) to get
${\mathbf{P}}^{N\times J}\left(\sup_{\psi_{z}\in\Psi}\left\|\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\psi_{z}(u^{i},y^{j})-{\mbox{\bf E}}_{(u,y)}[\psi_{z}(u,y)]\right\|\leq\tau\right)\geq 1-\left(\frac{D\sqrt{N}C\tau}{\sqrt{V}}\right)^{V}\exp(-2NC^{2}\tau^{2}).$
By the convergence of Algorithm 1 to an $\epsilon$-stationary point of (10), we have that the empirical $\mathop{\bf CVaR}$ will be close to $\kappa$, which, given equation (13), means
$\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\psi_{z}(u^{i},y^{j})\leq(1/C)(c^{\min}+\kappa+\epsilon).$
Therefore, we have
${\mathbf{P}}^{N\times J}\left(\sup_{\psi_{z}\in\Psi}{\mbox{\bf E}}_{(u,y)}[\psi_{z}(u,y)]\leq(1/C)(c^{\min}+\kappa+\epsilon)+\tau\right)\geq 1-\beta,$
where $\beta=(D\sqrt{N}C\tau/\sqrt{V})^{V}\exp(-2NC^{2}\tau^{2})$. If $\kappa$ is set such that $\kappa\leq-C\tau-\epsilon-c^{\min}$, we then have
${\mathbf{P}}^{N\times J}\left(\mathop{\bf CVaR}(g(x,u,y),\eta)\leq 0\quad\forall x\in\mathcal{X}\right)\geq 1-\beta,$
which implies
${\mathbf{P}}^{N\times J}\left({\mathbf{P}}_{(u,y)}(g(x,u,y)\leq 0)\geq 1-\eta\quad\forall x\in\mathcal{X}\right)\geq 1-\beta.$ (16)
As this probabilistic guarantee holds for all $x$ within the bounded set $\mathcal{X}$, it also holds for our data-driven solutions $x(\theta,y)$.
Corollary 4.1.1.
Suppose Assumption 3.4 holds. For all $\theta^{0}\in\mathcal{D}_{\theta}$, $s^{0}\in\mathcal{D}_{s}$, and $y\in\mathcal{D}_{y}$, Theorem 4.1 implies the probabilistic guarantee (7) for the data-driven solution $x(\theta^{\star},y)$, i.e., ${\mathbf{P}}^{N\times J}\left\{{\mathbf{P}}_{(u,y)}(g(x(\theta^{\star},y),u,y)\leq 0)\geq 1-\eta\right\}\geq 1-\beta.$
We note that the covering number requirement holds when $\Psi$ is a class of Lipschitz functions, which is implied when the constraint function $g$ is Lipschitz. For instance, when $g$ is affine or a maximum of affine functions in variable $x$, it is Lipschitz in $x$ for all $u\in\mathcal{D}_{u}$ and $y\in\mathcal{D}_{y}$,  i.e.,
$|g(x,u,y)-g(x^{\prime},u,y)|\leq\bar{L}\|x-x^{\prime}\|_{2}\quad\forall x,x^{\prime}\in\mathcal{X},~{}u\in\mathcal{D}_{u},~{}y\in\mathcal{D}_{y},$
with some Lipschitz constant $\bar{L}$. Then, $\psi_{z}\in\Psi$ is Lipschitz in $z=(x,\alpha)$ with Lipschitz constant $L\leq(1/C)((\bar{L}+1)/\eta+1)$. We can now derive the following covering number condition.
Covering number for $L$-Lipschitz functions.
When $\psi_{z}$ is $L$-Lipschitz in $z$ for all $u$ and $y$, we can apply standard bounds on the covering number of Lipschitz losses (van der Vaart and Wellner,, 1996; Pollard,, 1890; Namkoong and Duchi,, 2017). In this case, the covering number condition (14) reduces to
$\sup_{Q}N(\delta,\Psi,L_{2}(Q))\leq N(\delta/L,\mathcal{Z},\|\cdot\|_{2})\leq\left(1+\frac{(x^{\rm max}+\alpha^{\rm max})L}{\delta}\right)^{n+1}\leq\left(\frac{K}{\delta}\right)^{n+1},$
for every $0<\delta<K$, where $K=\delta+(x^{\rm max}+\alpha^{\rm max})L$. Then, $\beta=(D\sqrt{N}C\tau/\sqrt{n+1})^{n+1}\exp(-2NC^{2}\tau^{2})$.
5 Examples
We run the experiments on the Princeton Institute for Computational Science and Engineering (PICSciE) facility with 20 parallel 2.4 GHz Skylake cores. We solve the inner optimization problems (1) with the Clarabel optimizer with default settings. The learning procedure relies on the Python package cvxpylayers (Agrawal et al., 2019a, ), and the code to replicate the experiments can be accessed at
https://github.com/stellatogrp/lropt_experiments
For each experiment, we compare our method with a RO and a DRO benchmark, given below.
Learning for robust optimization (LRO) (our method).
We apply the training method given in Section 3.2, for a variety of $\eta$ values, to achieve various empirical probabilities of constraint violation $\hat{\eta}$. The uncertainty set can be RO or DRO based. Default training parameters are given in Table 2.
LRO + Fine Tuning (our method).
Upon completion of the LRO training procedure, we also fine tune the size of the uncertainty set using grid search to precisely achieve a target in-sample empirical probability of constraint violation.
Mean-Variance RO.
We perform mean-variance RO, where the shape of the uncertainty set is determined by the empirical mean and variance of the data (Ben-Tal and Nemirovski,, 2000). We use grid search over a the set size to achieve various empirical probabilities of constraint violation $\hat{\eta}$.
Wasserstein DRO.
We perform Wasserstein-1 DRO (Esfahani and Kuhn,, 2018), where the uncertainty set is constructed around individual data points. We again use grid search over the set size to achieve various empirical probabilities of constraint violation $\hat{\eta}$.
For each setting, we repeat the experiment 20 times, each with a new independent dataset. For each method, we evaluate the performance metrics in Table 1.
Table 1: Performance metrics in numerical examples, averaged over 20 runs.
Name Description Obj. Out-of-sample objective value estimate for problem (8) $\hat{\eta}$ Out-of-sample empirical probability of constraint violation $\hat{\beta}$ Out-of-sample estimated confidence, given as ${\mathbf{P}}(\hat{\eta}>\eta)$, where the probability is over the testing data ${\rm time}$ Average solution time of the final robust problem after training
Table 2: Default hyperparameter values for Algorithm 1. We set the initial uncertainty set parameter values $(A^{0},b^{0})=(\hat{\Sigma}^{1/2},\hat{\mu})$, depending on the empirical covariance and mean of the training data.
$\theta^{0}$ $\alpha^{0}$ $s^{0}$ $\lambda^{0}$ $\mu^{0}$ $\kappa$ $\sigma$ $\tau$ $k_{\text{max}}$ $t_{\text{max}}$ $\eta$ $|B|$
$(A^{0},b^{0})$ 0 0 1 1 $-0.01$ 1.015 0.95 1000 100 $[0,0.3]$ 30
5.1 Portfolio optimization
Problem description.
We consider the portfolio management problem where we select a portfolio $x\in{\mbox{\bf R}}^{n}$ of stocks to maximize $u^{T}x$, where $u\in{\mbox{\bf R}}^{n}$ are the uncertain asset returns. The trading costs are proportional to the deviations from our nominal holdings $x^{\text{nom}}\in{\mbox{\bf R}}^{n}$, i.e., $\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}$ with $\lambda^{\rm trade}>0$. We seek to minimize the objective $-u^{T}x+\lambda^{\rm trade}\|x-x^{\text{nom}}\|_{1}.$ We parametralize the problem with the family parameter $y=x^{\text{nom}}$, and solve the RO problem
$\begin{array}[]{ll}\mbox{minimize}&t+\lambda^{\text{trade}}\|x-x^{\text{nom}}\|_{1}\\ \mbox{subject to}&-u^{T}x\leq t\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}x=1,\quad x\geq 0,\end{array}$
where the uncertain parameter $u$ stands for the unknown returns.
Data.
We consider the case where $n=5$ and $n=10$ stocks. We let $\lambda^{\text{trade}}=0.2$. The uncertain returns $u$ is sampled from a multivariate normal distribution $N(\mu,\Sigma)$, where $\mu$ is generated uniformly on $[0.5,1]^{n}$, and $\Sigma$ is generated as $0.1FF^{T}$, where $F$ is generated on $[0,1]^{n\times 2}$. The nominal holdings $\bar{x}^{\text{nom}}$ is generated from a Dirichlet distribution with parameter values equal to $\mu$. We consider 10 instances of $x^{\text{nom}}$, distributed normally around $\bar{x}^{\text{nom}}$, with a variance of $0.05$. Each sample is then projected to be non-negative, and normalized to sum to 1. The total training sample size is $N=500$, with a testing sample size of $20000$.
Uncertainty set.
We learn both a RO and a DRO uncertainty set. For RO, the set is ellipsoidal with with norm $p=2$, and for DRO, we train the shape of a mean robust optimization (MRO) set with $K=5$ and power $p=1$. This set is closely related to the Wasserstein type-1 DRO set. See Appendix A for details on the MRO uncertainty set. By choosing $K\ll N$, we speed up the training procedure with minimal performance degradation compared to training with Wasserstein DRO (Wang et al.,, 2022). For both $n=5$ and $n=10$, we learn uncertainty sets for $\eta\in\eta^{\text{list}}=(0.01,0.03,0.05,0.08,0.10,0.15,0.2,0.3)$.
Results.
In Table 3, we compare the performance of our LRO methods with mean-variance RO and Wasserstein DRO, and note the desirable trade-off given by our learned sets.
[…]
In Figure 3, for the 20 runs of the experiment, and for $m\in\{5,10\}$, we show the trade-off between the average objective value and empirical probability of constraint violation. The trade-offs for the LRO sets are obtained by applying the training algorithm for the $\eta$ values provided above, fine-tuning the size of the uncertainty set, and reporting the best objective values for a range of $\hat{\eta}$ values. We observe lower objective values for the LRO sets across the relevant levels of constraint violation. Note that in general, at the same out-of-sample objective value, the mean-variance RO and nominal Wasserstein DRO sets suffer from a much higher empirical probability of constraint violation. Therefore, the decision-focused uncertainty sets are more effective than uncertainty sets based on the distribution alone.
Comparing the LRO-RO and LRO-MRO sets, we note that the RO-based set has a higher variance in performance. However, as seen for $n=10$, it could outperform the MRO-based set. In terms of computation time, while the RO problems are less expensive to solve, the MRO framework (Wang et al.,, 2022) already presents a significant computational time reduction compared to Wasserstein DRO, especially when $K\ll N$.
In Figure 4, for $n=10$ stocks, a single $y$, and a single run of the experiment, we show the distribution of the optimal portfolio weights for all four uncertainty sets. The mean of the uncertain returns is $\mu=(0.5,0.55,0.57,0.59,0.65,0.67,0.7,0.71,0.77,0.86)$, sorted in ascending order (bottom to top in Figure 4), and the corresponding diagonals of $\Sigma$ are $(0.64,0.08,0.25,0.03,0.06,0.25,0.11,0.36,0.13,0.02)$. The covariances between the entries are omitted here for brevity. To mitigate the risks, when $\hat{\eta}$ is low the distribution of portfolio weights is either skewed towards the relatively lower variance stocks, as is the case for the the RO sets, or is close to an even distribution, as is the case for the DRO sets. Intuitively, one would invest more in stocks with higher returns and higher risks when the proxy of risk tolerance, $\hat{\eta}$, is high. Indeed, we observe that the relatively high-risk, high-return stock, the third-to-last stock (depicted in orange), receives larger investments as $\hat{\eta}$ increases. Note, however, that the reshaped sets, LRO-RO and LRO-MRO, are able to identify this profitable stock and increase its portfolio weight even at relatively low values of $\hat{\eta}$. By identifying a more compact region of uncertainty targeted for the problem family, they are able reduce conservatism in the portfolio selection process and take more profitable positions.
Figure 3: Portfolio. Average trade-offs between out-of-sample objective values and probability of constraint violation across 20 experiments. The shaded region represents the 0.3 to 0.7 quantiles. The LRO results are obtained through training procedures with $\eta$ values in the range $[0,0.3]$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Portfolio. Optimal portfolio weights chosen by the four methods for various $\hat{\eta}$. The colors represent the true mean $\mu$ of the returns.
5.2 Inventory management
Problem description.
We consider a two-stage adjustable RO problem on inventory management (Iancu and Trichakis,, 2014, Section 5.2.1). There exist a retail network consisting of a single warehouse and $n$ different retail points, indexed by $i=1,\dots,n$. For simplicity, only a single product is sold. There are a total of $C$ units of the product available, and each retail point starts with 00 inventory and is capable of stocking at most $c_{i}$ units. The transportation costs for distributing the items at the $i$th retail point is $t_{i}$ currency units per unit of inventory, and the operating cost is $h_{i}$ per unit. The revenues per unit is $r_{i}$, which we take as our family parameter, ${\it i.e.},~{}y=r$, as these values are known but may be subject to minor changes across problem instances. Customer demand is uncertain, with a nominal demand plus an additional amount dependent on $m$ market factors. The formula for the demand at retail point $i$ is
$a_{i}=\bar{a}_{i}+q_{i}^{T}u,$
where $\bar{a}_{i}\in{\mbox{\bf R}}$ denotes the nominal demand, and $q_{i}\in{\mbox{\bf R}}^{m}$ denotes the exposure of demand $a_{i}$ to the market factors. We denote with $u$ the unknown market factors, which belong to an uncertainty set ${\mathcal{U}}(\theta)$. The decision variable is the stocking decisions for all retail points, denoted $s\in{\mbox{\bf R}}^{n}$. In addition, the problem involves a vector $w(u)$ of realized sales at each point, which is dependent on stocking decision $s$ and the unknown factors $u$ affecting the demand. We have
$w_{i}(u)=\min\{s_{i},a_{i}\},$
for each retail point. The manager makes stocking decisions to maximize worst-case profits, which is equivalent to minimizing worst-case loss, $-r^{T}w(u)+(t+h)^{T}s$, $\forall u\in{\mathcal{U}}(\theta)$. For each problem instance, the RO formulation is given as
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}w(u)+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta)\\ &w_{i}(u)\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &w_{i}(u)\leq\bar{a}_{i}+q_{i}^{T}u,\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (17)
where variables $s\in{\mbox{\bf R}}^{n}$, $w\in{\mbox{\bf R}}^{n}$, and $\tau\in{\mbox{\bf R}}$. To get a tractable approximation, we replace the adjustable decision $w_{i}(u)$ with their affine adjustable robust counterparts $z^{0}_{i}+v_{i}^{T}u$, where $z^{0}_{i}\in{\mbox{\bf R}}$ and $v_{i}\in{\mbox{\bf R}}^{f}$ are auxiliary introduced as part of the linear decision rules (Ben-Tal et al.,, 2009). The new formulation is
$\begin{array}[]{ll}\mbox{minimize}&\tau\\ \mbox{subject to}&-r^{T}z^{0}-r^{T}Vu+(t+h)^{T}s\leq\tau,\quad\forall u\in{\mathcal{U}}(\theta)\\ &z^{0}_{i}+v_{i}^{T}u\leq s_{i},\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &z^{0}_{i}+v_{i}^{T}u\leq\bar{a}_{i}+q_{i}^{T}u,\quad i=1,\dots,n,\quad\forall u\in{\mathcal{U}}(\theta)\\ &\mathbf{1}^{T}s=C,\quad 0\leq s\leq c,\\ \end{array}$ (18)
where $V=[v_{1}\cdots v_{n}]^{T}\in{\mbox{\bf R}}^{n\times f}$. We treat the multiple uncertain constraints as a single joint chance constraint represented as the maximum of concave functions.
Data.
As in the work by Iancu and Trichakis, (2014), we let $n=10$. We consider the cases where the number of uncertain market factors is $m=4$ and $m=8$. We set $C=200$, and all other problem data are independently sampled from uniformly distributed random variables. Inventory capacities $c$ are between 30 and 50. Transportation and operation costs are between 1 and 3. Sales revenues $r$ are between 2 and 4, and nominal demand values are between 10 and 20. We consider 10 instances for $r$, perturbing each by random sample of a normal distribution with mean 0 and standard deviation $0.05$. The market factor exposure coefficients $q_{i}$ are between $-2$ and $2$. We generate uncertain market factors $u$ from 3 normal distributions, to simulate 3 uncertain market environments. For distribution $j$, the mean is $\mu_{i}=i\gamma_{j}$ for all $i=1,\dots,m$, with $\gamma=(0.3,0.6,0.9)$, and the variance is $\sigma_{i}=0.02^{2}+(0.1i)^{2}$ for all distributions. The total training sample size is $N=100$, split evenly across the 3 distributions. The testing sample size is $20000$.
Uncertainty set.
We train the shape and size of an ellipsoidal uncertainty set, with norm $p=2$, and for $\eta\in\eta^{\text{list}}=(0.01,0.03,0.05,0.08,0.10,0.15,0.2,0.3)$.
Results.
In Table 4, we compare the performance of our LRO method with mean-variance RO and Wasserstein DRO, and note the desirable trade-off given by our learned reshaped set.
Table 4: Average inventory management performance over 20 runs. LRO: learning the shape and size of the uncertainty set with $\eta=0.3$. LRO-T0.03: Best value given by LRO with $\eta\in\eta^{\text{list}}$, tuned for an in-sample empirical probability of constraint violation ($\hat{\eta}^{\text{in}}$) below 0.03. LRO-T0.05: Best value given by LRO with $\eta\in\eta^{\text{list}}$, tuned for an $\hat{\eta}^{\text{in}}$ below 0.05. Mean-Var RO0.03: mean-variance RO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.03$. Mean-Var RO0.05: mean-variance RO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.05$. W-DRO0.03: Wasserstein DRO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.03$. W-DRO0.05: Wasserstein DRO where the size is tuned for an $\hat{\eta}^{\text{in}}$ below $0.05$. Obj.: the average out-of-sample objective value ${\mbox{\bf E}}[-r^{T}w(u)+(t+h)^{T}s]$. ${\rm time}$: average solve time after training.
[…]
In Figure 5, we show the trade-off between the average objective value and empirical probability of constraint violation for the 20 runs of the experiment. We observe an improved objective value for the LRO set across relevant levels of constraint violation. In Figure 5, for $m=4$, we note that the LRO set with a constraint violation $\hat{\eta}$ of around 0.028 achieves an objective value that the mean-variance RO and Wasserstein DRO sets only achieve at signficantly higher $\hat{\eta}$’s, of around 0.12 and 0.07, respectively. A similar comparison holds for $m=8$. It is apparent that our decision-focused manner of learning uncertainty sets identifies more effective regions of uncertainty and therefore reduces the conservatism of the data-driven solutions.
Figure 5: Inventory. Average trade-offs between out-of-sample objective values and probability of constraint violation across 20 experiments. The shaded region represents the 0.3 to 0.7 quantiles. The LROPT results are obtained through training procedures with $\eta$ values in the range $[0,0.3]$.
6 Conclusion
We develop a data-driven method to learn the shape and size uncertainty sets in robust optimization in a decision-focused manner, i.e., using the structure and the solution of the robust counterpart. By formulating the learning problem as a stochastic optimization problem with $\mathop{\bf CVaR}$ constraints, we choose the shape and size of the uncertainty sets to optimize the data-driven performance while enforcing desired probabilistic guarantees. We show convergence of our algorithm using nonsmooth implicit differentiation, and provide finite sample probabilistic guarantees on unseen data. Through numerical examples, we demonstrate the performance gain achieved by our method compared to traditional manual tuning methods.
7 Acknowledgments
Irina Wang and Bartolomeo Stellato are supported by the NSF CAREER Award ECCS 2239771. The simulations presented in this article were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center and Visualization Laboratory at Princeton University.
Appendix A Uncertainty sets and general reformulations
We derive convex reformulations of (1) for a general class of constraints $g$ and common uncertainty sets. We express $u$ as an affine transformation of an uncertain parameter $z$, i.e., $u=Ax+b$, where $A\in{\mbox{\bf R}}^{m\times\hat{m}}$, and $\hat{m}\leq m$, and $b\in{\mbox{\bf R}}^{m}$. The proofs are in Appendix E. For notational simplicity, we suppress the dependency of $g$ on $y$ here and for all following sections on reformulations. We consider two general types of functions $g$: the maximum of concave functions (left), where each $g_{l}$ is concave in $u$, and its special case, the maximum of affine functions (right), where $P_{l}(x)$ and $a_{l}(x)$ are convex functions in $x$:
$g(x,u)=\max_{l=1,\dots,L}g_{l}(x,u),\qquad g(x,u)=\max_{l=1,\dots,L}(P_{l}(x)u+a_{l}(x)).$
Box and Ellipsoidal uncertainty.
The ellipsoidal uncertainty set is defined as
${\mathcal{U}}(\theta)=\left\{u=Az+b\mid\|z\|_{p}\leq 1,\quad u\in S\right\},$
where $A\in{\mbox{\bf R}}^{m\times\hat{m}}$, $b\in{\mbox{\bf R}}^{m}$, $p\in{\mbox{\bf R}}$ , and $S$ includes additional support information such as upper and lower bounds. We let $p$ be the order of the norm, which is any integer value $\geq 1$, or equal to $\infty$. In the latter case, this becomes the box uncertainty set. The trainable parameter $\theta=(A,b)$. Note that if $A=\rho I$, $b=0$, and $S={\mbox{\bf R}}^{m}$, this is equivalent to the traditional ellipsoidal uncertainty set ${\{u\mid\|u\|_{p}\leq\rho\}}$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}\text{minimize}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|A^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L,\end{array}$
with auxiliary variables $\omega_{l},h_{l}\in{\mbox{\bf R}}^{m}$ for $l=1,\dots,L$, and $g^{\star}(x,h)$ the conjugate function. This result follows from the theorem of infimal convolutions and composites for conjugate functions (Rockafellar and Wets,, 1998)(Bauschke and Combettes,, 2017). As a special case, we have the reformulation for the maximum of affine constraint,
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega_{l})+\|A^{T}(\omega_{l}-P_{l}(x))\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
Budget uncertainty.
The budget uncertainty set requires both box and ellipsoidal constraints, and is defined as
${\mathcal{U}}_{\text{bud}}(\theta)=\{u=Az+b\mid\|z\|_{\infty}\leq\rho_{1},\quad\|z\|_{1}\leq\rho_{2},\quad u\in S\},$
where $\theta=(A,b,\rho_{1},\rho_{2})$. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L,\end{array}$
with $r_{l}\in{\mbox{\bf R}}^{\hat{m}}$, $\omega_{l},h_{l}\in{\mbox{\bf R}}^{m}$ for $l=1,\dots,L$ as auxiliary variables. The maximum of affine reformulation is
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+\sigma_{S}(\omega_{l})+b^{T}(P_{l}(x)-\omega_{l})+\rho_{1}\|r_{l}-A^{T}(P_{l}(x)+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\quad l=1,\dots,L.\end{array}$ (19)
Polyhedral uncertainty.
The polyhedral uncertainty set is defined as
${\mathcal{U}}(\theta)=\{u=Az+b\mid Dz\leq d,\quad u\in S\},$
where $\theta=(D,d)\in{\mbox{\bf R}}^{r\times\hat{m}}\times{\mbox{\bf R}}^{r}$ defines the polyhedron. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
with auxiliary variables $\gamma_{l},\omega_{l}\in{\mbox{\bf R}}^{m}$, $h_{l}\in{\mbox{\bf R}}^{r},$ for $l=1,\dots,L$. For the maximum of affine constraint, we have:
$\begin{array}[]{ll}\mbox{minimize}&f(x)\\ \mbox{subject to}&a_{l}(x)+d^{T}h_{l}+b^{T}(P_{l}(x)-\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(P_{l}(x)-\gamma_{l})=D^{T}h_{l},\quad l=1,\dots,L\\ &h_{l}\geq 0,\quad l=1,\dots,L.\end{array}$
Mean robust uncertainty.
The Mean Robust Uncertainty (MRO) (Wang et al.,, 2022) set is defined as
${\mathcal{U}}_{\text{mro}}(\theta)$ =$=\left\{u=A(v_{11},\dots,v_{LK})+b\in S^{K\times L}\quad\middle|\quad\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}\|v_{lk}-\bar{d}_{k}\|^{p}\leq\rho^{p},\quad\alpha\in\Gamma\right\},$
where $\Gamma=\{\alpha~{}|~{}\sum_{l=1}^{L}\alpha_{lk}=w_{k},\alpha_{lk}\geq 0~{}\forall k,l\}$. Here, we assume the constraint function $g$ of the RO to be a maximum of $L$ pieces, indexed by $l$. We assume A is square, with dimensions ${\mbox{\bf R}}^{m\times m}$, and we have $\theta=(A,b)$. This set depends on $U^{N}$, a dataset of $N$ samples. We partition ${U}_{N}$ into $K$ disjoint subsets, with $\bar{d}_{k}$ the centroid of the $k$th subset. The weight we place upon each subset is $w_{k}$, and is equivalent to the proportion of points in the subset. We choose $p$ to be an integer exponent, and $\rho$ to be a positive radius. For the maximum of concave constraint, we have the following reformulation.
$\begin{array}[]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$
with auxiliary variables $s,h_{lk},\omega_{lk}\in{\mbox{\bf R}}^{m}$ for $j=1,\dots,J$, $k=1,\dots,K$, and $\lambda\in{\mbox{\bf R}}$. For the special case of the maximum of affine constraint, we have
minimize $\begin{array}[]{ll}\text{minimize }&f(x)\\ \mbox{subject to}&\lambda\rho^{p}+\sum_{k=1}^{K}w_{k}s_{k}\leq 0\\ &a_{l}(x)+\sigma_{S}(\omega_{lk})-b^{T}(\omega_{lk}-P_{l}(x))-(A^{T}(\omega_{lk}-P_{l}(x)))^{T}\bar{d}_{k}\\ &\qquad\qquad+\phi(r)\lambda\left\|A^{T}(\omega_{lk}-P_{l}(x))/\lambda\right\|^{r}_{q}\leq s_{k},\quad l=1,\dots,L,\quad k=1,\dots,K\\ &\lambda\geq 0.\end{array}$ (20)
A.1 Linear transformation to cone programs in standard form
The convex optimization problems above can be transformed to cone programs (Agrawal et al., 2019a, ) in the form
$\begin{array}[]{ll}\mbox{minimize}&c^{T}x\\ \mbox{subject to}&Gx+s=h\\ &s\in\mathcal{K},\end{array}$ (21)
where $x\in{\mbox{\bf R}}^{\bar{n}}$ is the primal variable, $s\in{\mbox{\bf R}}^{\bar{m}}$ is a slack variable, and the set $\mathcal{K}\subseteq{\mbox{\bf R}}^{\bar{m}}$ is a nonempty, closed, convex cone with dual cone $\mathcal{K}^{\star}\subseteq{\mbox{\bf R}}^{\bar{m}}$. So long as the constituent functions $g_{l}$ of the constraint $g$ from the robust problem (1) are concave in the uncertain parameters, and that these functions satisfy the (disciplined parametrized program) DPP rules proposed by Agrawal et al., 2019a , the cone program’s data $(G,h,c)$ can be expressed as a linear function (i.e., tensor product) of the parameters $\theta$ and $y$ (Agrawal et al., 2019a, , Lemma 1). The DPP rules depend on well-known composition theorems for convex functions, and ensure that every function in the program is convex in the decision variables, or affine if representing equalities.
Appendix B Conservative Jacobians for the augmented Lagrangian
We give the exact forms of the conservative Jacobians from Proposition 3.1. In this section, for simplicity, we denote the solution of the optimization problem (1), $x^{j}=x(\theta,y^{j})$, and its conservative Jacobian, $J_{x^{j}}(\theta)$, for each input pair $(\theta,y^{j})$. For the Lagrangian function $\hat{L}(\alpha,s,\theta,\lambda,\mu)$ over the batch $B=(U\times Y)\subseteq(U^{N}\times Y^{J})$, where $|B|$ denotes the cardinality of the set, we have the following conservative Jacobians.
$J_{\hat{L}}(\theta)$ $=\frac{1}{|Y|}\sum_{y^{j}\in Y}J_{f}(x^{j})J_{x^{j}}(\theta)$
$+\frac{\lambda}{\eta|B|}\sum_{(u^{i},y^{j})\in B}\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\}J_{g}(x^{j})J_{x^{j}}(\theta)$
$+\frac{\mu}{\eta|B|}\sum_{(u^{i},y^{j})\in B}((g(x^{j},u^{i},y^{j})-\alpha)/\eta+\alpha+s-\kappa)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\}J_{g}(x^{j})J_{x^{j}}(\theta),$
$J_{\hat{L}}(s)$ $=\lambda+\frac{\mu}{|B|}\sum_{(u^{i},y^{j})\in B}(g(x^{j},u^{i},y^{j})-\alpha)_{+}/\eta+\alpha+s-\kappa,$
$J_{\hat{L}}(\alpha)$ $=\frac{\lambda}{|B|}\sum_{(u^{i},y^{j})\in B}(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\})$
$+\frac{\mu}{|B|}\sum_{(u^{i},y^{j})\in B}(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha\right\})((g(x^{j},u^{i},y^{j})-\alpha)_{+}/\eta+\alpha+s-\kappa).$
For a condition $A$, we denote its indicator function $\mathbf{1}\{A\}$, taking the value 1 when true and 0 when false. When the Lagrangian function is defined over the entire dataset, we formulate the conservative Jabocian $J_{L}(\alpha,s,\theta)$ the same way as above, with $B=(U^{N}\times Y^{J})$.
Appendix C Proof of conservative Jacobians for the robust optimization problem
To prove Theorem 3.1, we characterize the path-differentiability of the function $x(\theta,y)$ by solving the inner optimization problem and obtaining conservative Jacobians using the results of Agrawal et al., 2019b and Bolte et al., (2021). The optimization problem (1) is first reformulated into a convex and tractable problem in conic form (21), as shown in Appendix A.1. The conic program’s data, $(G,h,c)$, is parameterized by $(\theta,y)$, through the mapping
$(\theta,y)\rightarrow(G,h,c),$
which is linear in $\theta$ and $y$ for all uncertainty sets in Appendix A, given the disciplined parametrized program (DPP) requirement detailed in Appendix A.1 (Agrawal et al., 2019a, , Lemma 1).
This formulation has a corresponding dual, with dual variable $r\in{\mbox{\bf R}}^{m}$, and solving the conic program amounts to solving the homogeneous self-dual embedding (Ye et al.,, 1994) of the primal-dual problem. In the following, we define $N=\bar{n}+\bar{m}+1$. The overall procedure is split into three steps: mapping the problem data ($G,h,c$) onto a skew symmetric matrix $Q$:
$Q=\left[\begin{array}[]{lll}0&G^{T}&c\\ -G&0&h\\ -c^{T}&-h^{T}&0\end{array}\right]\in{\mbox{\bf R}}^{N\times N},$
finding the solution $z=(u,v,w)\in{\mbox{\bf R}}^{N}$ to the homogeneous self-dual embedding, and mapping $z$ onto the solution of the primal-dual pair (Agrawal et al., 2019b, ). We can express the chain of mappings as
$(G,h,c)\rightarrow Q\rightarrow z\rightarrow(x,r,s).$
To solve for $z=\nu(Q)$, shown as the mapping $Q\rightarrow z$, we define the normalized residual map (Busseti et al.,, 2019), given as
$\mathcal{N}(z,Q)=((Q-I)\Pi+I)(z/|w|),$
where $\Pi$ is the projection onto the convex cone $\mathcal{K}^{\prime}={\mbox{\bf R}}^{\bar{n}}\times\mathcal{K}^{\star}\times{\mbox{\bf R}}_{+}$. For a given $Q$, $z=\nu(Q)$ is a solu tion to the homogeneous self-dual embedding if and only if $\mathcal{N}(z,Q)=0$ and $w>0$. Once this solution is obtained, the mapping $z\rightarrow(x,r,s)$ is given by the function
$\phi(z)=(u,\Pi(v),\Pi(v)-v)/w.$
We can now invoke proposition 4 of the work by Bolte et al., (2021), together with the results from Agrawal et al., 2019b to show that $\nu$ and $\phi$ are path-differentiable.
Proposition C.1.
Assume the functions $\Pi$ and $\mathcal{N}$ are path differentiable, with conservative Jacobians $J_{\Pi}$ and $J_{\mathcal{N}}$. Assume, for all $Q$ constructed from the problem data $(G,h,c)$, and $z=\nu(Q)$ is the unique solution to $\mathcal{N}(z,Q)=0$. Furthermore, assume all matrices formed by the first $N$ columns of $J_{\mathcal{N}}(z,Q)$ are invertible. Then, the functions $\nu$ and $\phi$ are path-differentiable, with conservative Jacobians
$J_{\nu}(Q)=\{-U^{-1}V\mid[U~{}V]\in J_{\mathcal{N}}(z,Q)\},\qquad J_{\phi}(z)=\left[\begin{array}[]{ccc}I&0&-x\\ 0&J_{\Pi}(v)&-r\\ 0&J_{\Pi}(v)-I&-s\end{array}\right].$
The assumptions on $\Pi$ and $\mathcal{N}$ hold, as the cones we consider (zero cones, free cones, nonnegative cones, and second-order cones) are sub-differentiable (Busseti et al.,, 2019; Kanzow et al.,, 2009). By Theorem 1 of the work by Bolte et al., (2021), which states that the Clarke Jacobian is a minimal conservative Jacobian, we conclude that the conic projections admit conservative Jacobians, and are thus path differentiable. It follows that $\mathcal{N}$, as a linear function of $\Pi$ and $Q$, is also path-differentiable. The invertibility of $U$ holds as $U=((Q-I)J_{\Pi}(z)+I)/w$ (Agrawal et al., 2019b, ), which is nonsingular due to the PSD nature of the members of the set $J_{\Pi}(z)$. Lastly, the assumption of uniqueness holds from Assumption 3.3.
With this result, and by the linear nature of the mappings from $\theta$ to $(G,h,c)$ and to $Q$, which preserves path-differentiablity, we can apply the chain rule to obtain
$J_{\phi}(\theta)=J_{\phi}(v(Q))J_{v}(Q)J_{Q}(G,h,c)J_{(G,h,c)}(\theta).$
As $(x,r,s)=\phi(z)$, we can partition $J_{\phi}(\theta)$ as $(J_{x}(\theta),J_{r}(\theta),J_{s}(\theta))$ where the first component is the conservative Jacobian of interest. Therefore, we conclude that $x(\theta,y)$ is path differentiable, with nonempty conservative Jacobian $J_{x}(\theta)$.
Appendix D Proof of convergence of the augmented Lagrangian algorithm
We prove Theorem 3.2, the convergence of the augmented Lagrangian algorithm to a stationary point. We separate the procedure into three steps. In the first step, we show for Algorithm 2 with fixed $\lambda,\mu$, the convergence of $\{\alpha^{t},s^{t},\theta^{t}\}$ to a stationary point of the augmented Lagrangian function (11) as $t\to\infty$. In particular, we note that by Theorem 1 of Bolte and Pauwels, (2019), conservative Jacobians are gradients almost everywhere, so by Bolte et al., (2021), the iterates of the algorithm is a Clarke stochastic subgradient sequence almost surely. We thus use the Clarke differential (Clarke,, 1975), $\partial^{c}f(x)\coloneqq{\rm conv}\{\lim_{x_{i}\rightarrow x}\nabla f(x_{i})\mid x_{i}\in{\rm dom}(\nabla f)\}$ to denote the subdifferentials in our analysis, instead of using $J_{f}(x)$. In the next two steps, we show for Algorithm 1, either the convergence of $\{\theta^{k}\}$ to a feasible point of conditon (9), or the convergence of $\{\alpha^{k},s^{k},\theta^{k}\}$ to a stationary point of the constraint violation function.
Step 1: Convergence of ($\alpha$, $s$, $\theta$).
In this step, we consider the augmented Lagrangian function with fixed $\lambda$ and $\mu$. By Assumptions 3.4 and 3.5 on the boundedness and step sizes of the iterates, by the definability of the augmented Lagrangian function, and by the existence of its conservative Jacobians, we apply Theorem 3 of Bolte et al., (2021), on the convergence of SGD for path-differentiable functions. The projection onto the nonnegative orthant for $s$ does not affect the convergence results, as proximal SGD with conservative Jacobians also converges, subject to the same assumptions (Davis et al.,, 2018, Theorem 6.2). Let $L(\alpha,s,\theta)=L(\alpha,s,\theta,\lambda,\mu)$, as the suppressed variables are fixed, and let $(\alpha^{\star},s^{\star},\theta^{\star})$ be a local minimum point of this function. By the aforementioned theorems, we can show that $V_{\lambda,\mu}(\alpha,s,\theta)\coloneqq L(\alpha,s,\theta)-L(\alpha^{\star},s^{\star},\theta^{\star})$ is a Lyapunov function for the differential inclusion $(\dot{\alpha},\dot{s},\dot{\theta})\in-\partial^{c}L(\alpha,s,\theta)$. Specifically, there exist a ball of radius $r$ around $(\alpha^{\star},s^{\star},\theta^{\star})$ such that for $(\alpha,s,\theta)\in\mathbf{B}_{\alpha^{\star},s^{\star},\theta^{\star}}(r)$, the Lyapunov function is locally positive definite. This function is decreasing along solutions strictly outside of the set of Clarke-critical points $\{(\alpha,s,\theta)\in{\mbox{\bf R}}^{p+2},0\in\partial^{c}L(\alpha,s,\theta)\}$. Then, for almost all initializations $(\alpha^{0},s^{0},\theta^{0})\in\mathcal{D}_{\alpha}\times\mathcal{D}_{s}\times\mathcal{D}_{\theta}$, the objective $L(\alpha^{t},s^{t},\theta^{t})$ converges almost surely and all stationary points $(\bar{\alpha},\bar{s},\bar{\theta})$ of $\{\alpha^{t},s^{t},\theta^{t}\}$ are Clarke-critical in the sense that $0\in\partial^{c}L(\bar{\alpha},\bar{s},\bar{\theta})$ (Bolte et al.,, 2021). We thus conclude that, with almost any initial point $(\alpha^{0},s^{0},\theta^{0})\in\mathbf{B}_{\alpha^{\star},s^{\star},\theta^{\star}}(r)\subseteq\mathcal{D}_{\alpha}\times\mathcal{D}_{s}\times\mathcal{D}_{\theta}$, the sequence $\{\alpha^{t},s^{t},\theta^{t}\}$ converges to a stationary point $(\alpha^{\star},s^{\star},\theta^{\star})$, satisfying $0\in J_{L}(\alpha^{\star},s^{\star},\theta^{\star})$.
Step 2: Convergence of $\lambda$.
From step 1, we have a stationary point $(\alpha^{k},s^{k},\theta^{k})$ of the augmented Lagrangian function at iteration $k$ of Algorithm 1, with fixed multiplier and penalty parameters $\lambda^{k-1}$ and $\mu^{k-1}$. Further properties of this stationary point depend on the convergence, or lack thereof, of the sequence $\{\lambda^{k}\}$. As the step size parameter of $\lambda$ is the penalty $\mu$, we consider the two cases, (a) where the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and (b) where $\{\mu^{k}\}$ is unbounded. We first analyze case (a). By lines 6-12 of Algorithm 1, we note that for $\{\mu^{k}\}$ to be bounded, the penalty parameter must remain constant from some iteration onwards. By the condition on line 6, with $\tau<1$, this implies $\lim_{k\rightarrow\infty}\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}=0$, which in turn implies $\lim_{k\rightarrow\infty}\mu^{k}(H(\alpha^{k},\theta^{k})+s^{k}-\kappa)=0$. This sequence of updates for $\lambda$ is geometric and summable, so the sequence $\{\lambda^{k}\}$ is bounded, with a convergent subsequence. Given the boundedness assumption on the true Lagrangian multiplier $\lambda^{\star}$ in Assumption 3.4, we conclude that the whole sequence converges, ${\it i.e.},$$\{\lambda^{k}\}\rightarrow\lambda^{\star}$. We now turn to case (b), where $\{\mu^{k}\}$ is unbounded. In this case, $\{\lambda^{k}\}$ may not converge. However, by the $\lambda$ update rule in lines 7 of Algorithm 1, we note that while $\{\mu^{k}\}\rightarrow\infty$, $\{\lambda^{k}\}$ is bounded between $\lambda^{\rm{min}}$ and $\lambda^{\rm{max}}$, the safeguard values.
Step 3: convergence of the outer loop.
We now show that either the sequence $\{\theta^{k}\}$ converges to a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition, or $\{\alpha^{k},s^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function. We begin with the former. Suppose the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and converges to $\mu^{\star}$. Then, as $k\to\infty$, $\{\lambda^{k}\}$ converges to the stationary point $\lambda^{\star}$. From step 1, we have that $\{\theta^{t}\}$ in the inner loop converges almost surely to $\theta^{k}=\theta^{\star}(\lambda^{k-1},\mu^{k-1})$ along each step of $\lambda$ and $\mu$’s convergence. In order to establish the convergence of $\{\theta^{k}\}$ in the outer loop to a stationary point $\theta^{\star}(\lambda^{\star},\mu^{\star})$, a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition (9), we must ensure that $\theta^{\star}(\lambda,\mu)$ is well-behaved, by showing its continuity with respect to $(\lambda,\mu)$.
Recall Assumption 3.3 on the uniqueness of the inner RO solutions $x(\theta,y)$, Assumption 3.4 on the compactness of the set $\mathcal{X}$, and the convexity (and therefore continuity) of the objective and constraint functions $f$ and $g$ in $x$. Suppose these assumptions hold. By Theorem 4.2.1 of Bank et al., (1984), we establish the continuity of $L(\alpha,s,\theta,\lambda,\mu)$ in $\theta$, for $\theta\in\mathcal{D}_{\theta}$. As $L$ is bilinear and therefore continuous in $(\lambda,\mu)$, we obtain our desired stability results, i.e., as $\{\lambda^{k}\}$ and $\{\mu^{k}\}$ converge to $\lambda^{\star}$, $\mu^{\star}$, the optimal solution with respect to $\theta$ is continuous, and converges almost surely to the local minimizer $\theta^{\star}$. In addition, by the strong convexity of $L$ in $s$, the optimal solution with respect to $s$ also converges to the global minimizer $s^{\star}$. Note that $\{\alpha^{k}\}$ does not necessarily converge to a stationary point, as the minimum over $\alpha$ is not necessarily unique (Uryasev and Rockafellar,, 2001, Theorem 2). However, by the convergence of the inner loop in step 1 and by the convergence condition of $\mu$ in step 2, as $k\to\infty$ in Algorithm 1, we have a point $(\alpha^{k},s^{\star},\theta^{\star},\lambda^{\star},\mu^{\star})$ satisfying $0\in J_{L}(\alpha^{k},s^{\star},\theta^{\star})$ and $H(\alpha^{k},\theta^{\star})+s^{\star}-\kappa=0$. We can conclude that $\{\theta^{k}\}$ converges almost surely to a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition, where the minimizer over $\alpha$ takes some value $\alpha^{\star}\leq\alpha^{k}$.
On the other hand, suppose the sequence of penalty parameters $\{\mu^{k}\}$ is unbounded. By step 1, for any $\lambda^{k}$ and $\mu^{k}$, we have obtained a stationary point of the augmented Lagrangian function, satisfying $0\in J_{L}(\alpha^{k},s^{k},\theta^{k})$. Note that, with $\{\mu^{k}\}$ unbounded and $\{\lambda^{k}\}$ bounded, we observe $\lim_{k\rightarrow\infty}\lambda^{k}/\mu^{k}=0$. Therefore, dividing both sides of the conservative Jacobian by $\mu^{k}$ and taking the limit, we obtain:
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{\eta NJ}\sum_{y^{j}\in Y^{J}}\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha^{k}\right\}((g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa)J_{g}(x^{j})J_{x^{j}}(\theta^{k}),$
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{NJ}\sum_{(u^{i},y^{j})\in(U^{N}\times Y^{J})}(g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa,$
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{NJ}\sum_{(u^{i},y^{j})\in(U^{N}\times Y^{J})}\left(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha^{k}\right\}\right)((g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa).$
These are the stationary conditions for the constraint violation function $V(\alpha,s,\theta)=(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$, i.e., $0\in J_{V}(\alpha^{k},s^{k},\theta^{k})$. We thus converge to a stationary point of the constraint violation function, as desired.
Appendix E General reformulation proofs
E.1 Conjugate of a general function
We first derive the conjugate of general functions, which can then be used in the upcoming reformulations.
Linear function.
For a linear function $g(x,u)=(Pu+a)^{T}x$, we compute the conjugate as
$[-g]^{\star}(x,h)=\sup_{u}h^{T}u-(a+Pu)^{T}x=\begin{dcases}a^{T}x&\text{ if}\;h=-P^{T}x\\ \infty&\text{ otherwise}.\end{dcases}$
Sum of concave functions.
For a general function $g(x,u)=\sum_{j=1}^{J}g_{j}(u,x)$, where each $g_{j}$ is concave in $u$, by the theorem of infimal convolutions (Rockafellar and Wets,, 1998, Theorem 11.23(a), p. 493), we compute the conjugate as
$[-g]^{\star}(x,h)=\begin{cases}\underset{h_{j},\dots,h_{J}}{\inf}&\sum_{j=1}^{J}[-g_{j}]^{\star}(x,h_{j})\\ \mbox{subject to}&\sum_{j=1}^{J}h_{j}=h.\end{cases}$
Maximum of concave functions.
We also consider the maximum of concave functions,
$g(x,u)=\max_{l=1,\dots,L}g_{l}(x,u).$
In practice, these functions $g(x,u)$ are often affine functions. For all the reformulations below, we assume $g(x,u)$ to be of this maximum of concave form.
E.2 Box and Ellipsoidal reformulations
Rewriting the robust constraint as optimization problem, we have
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&\|z\|_{p}\leq 1\end{array}$
$=\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}g(x,Az+b)-\lambda\|z\|_{p}+\lambda\\ \end{array}$
[…]
$=\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-\lambda_{l}\|z\|_{p}\leq\tau,\quad l=1,\dots,L,\end{array}$
where $p=\infty$ for box uncertainty and $p\geq 1$ otherwise. We used the Lagrangian in the first equality, and decomposed $g(x,Az+b)$ into its constituents for the second equality. As $l$ and $\lambda$ are separable, in the third step we interchanged the $\inf$ and the $\max$. Next, for each $l$, as the expression is convex in $\lambda$ and concave in $u$, we used the Von Neumann-Fan minimax theorem to interchange the $\inf$ and the $\sup$. We then express the $\max$ in epigraph form by introducing an auxiliary variable $\tau$.
Now, if we define the function $c(z)=\|z\|_{p}$, then using the definition of conjugate functions, we have, for each $l$, the left-hand-side of the constraint equal to
$\lambda_{l}+[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0).$
We then borrow results from (Rockafellar and Wets,, 1998, Theorem 11.23(a), p. 493)(Zhen et al.,, 2021, Lemma C.9), and (Bauschke and Combettes,, 2017, Proposition 13.24) with regards to the inf-convolution of conjugate functions, the conjugate functions of norms, and the conjugate of linear compositions, to note
$[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0)=\underset{\omega_{l},h_{l},y_{l}}{\inf}([-g_{l}]^{\star}(x,h_{l})$ $-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|y_{l}\|_{q}\leq\lambda_{l}}$
$+\mathcal{I}_{A^{T}(h_{l}+\omega_{l})=y_{l}}),$
where $h_{l}\in{\mbox{\bf R}}^{m}$, $y_{l}\in{\mbox{\bf R}}^{\hat{m}}$, $\omega_{l}\in{\mbox{\bf R}}^{m}$, and $q$ is the conjugate number of $p$. Substituting these back, we arrive at
$\begin{array}[t]{ll}\underset{x,\tau,h,\omega,y}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|y_{l}\|_{q}\leq\tau,\quad l=1,\dots,L\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
Recall that this optimization problem upper bounds our original robust constraint, which we constrain to be nonpositive. Therefore, we only need $\tau\leq 0$. For the final step, we can now write the original minimization problem (1) in minimization form for both the objective and constraints, to arrive at
$\begin{array}[t]{ll}\underset{x,h,\omega}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|A^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
E.3 Budget reformulation
Rewriting the robust constraint as optimization problem, we have
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&\|z\|_{\infty}\leq\rho_{1}\\ &\|z\|_{1}\leq\rho_{2}\\ \end{array}$
[…]
where we again used the Lagrangian, separated $g$ into its constituents, and applied the minimax theorem. We now express the max in epigraph form,
$\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-\lambda_{l1}\|z\|_{\infty}-\lambda_{l2}\|z\|_{1}\leq\tau,\quad l=1,\dots,L.\end{array}$
If we define the functions $c_{1}(z)=\|z\|_{\infty}$, $c_{2}(z)=\|z\|_{1}$, then using the definition of conjugate functions, we have for each $l$ the left-hand-side of the constraint equal to
$\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(0)$
$=\underset{h_{l},\omega_{l},y_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(x,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+[\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(-y_{l})$
$\hskip 341.43306pt+\mathcal{I}_{A^{T}(h_{l}+\omega_{l})=y_{l}}$
$=\underset{h_{l},\omega_{l},y_{l},r_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(x,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|r_{l}-y_{l}\|_{1}\leq\lambda_{l1}}+\mathcal{I}_{\|r_{l}\|_{\infty}\leq\lambda_{l2}}$
$\hskip 341.43306pt+\mathcal{I}_{A^{T}(h_{l}+\omega_{l})=y_{l}}$
We can express this as
$\begin{array}[t]{ll}\underset{x,h_{l},\omega_{l},y_{l},r_{l},\tau}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-y_{l}\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq\tau,\\ &\hfill k=1,\dots,K\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
We can then let $\tau=0$ and substitute the new constraints in for the robust constraint in (1),
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\\ &\hfill l=1,\dots,L.\end{array}$
E.4 Polyhedral reformulation
Rewriting the robust constraint as an optimization problem, we have
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&Dz\leq d\end{array}$
$=\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}~{}\underset{h\geq 0}{\inf}~{}g(x,Az+b)-(Dz)^{T}h+d^{T}h\\ \end{array}$
[…]
$=\begin{array}[t]{ll}\underset{h\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-(Dz)^{T}h_{l}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{h\geq 0,\omega,\tau,y,\gamma}{\inf}&\tau\\ \mbox{subject to}&d^{T}h_{l}+[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})+\mathcal{I}_{D^{T}h_{l}=-y_{l}}\\ &\hfill+\mathcal{I}_{A^{T}(\omega_{l}+\gamma_{l})=y_{l}}\leq\tau,\quad l=1,\dots,L\end{array}$
$=\begin{array}[t]{ll}\underset{z\geq 0,\omega,\tau,\gamma}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq\tau,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L,\end{array}$
where we again used the Lagrangian, the Von Neumann-Fan minimax theorem, the infimal convolution of conjugates, and the conjugate of affine compositions. The final reformulation (1) is then
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,\omega_{l})-b^{T}(\omega_{l}+\gamma_{l})+\sigma_{S}(\gamma_{l})\leq 0,\quad l=1,\dots,L\\ &A^{T}(\omega_{l}+\gamma_{l})=-D^{T}h_{l},\quad l=1,\dots,L.\end{array}$
E.5 Mean Robust Optimization reformulation
We adapt the proof from (Wang et al.,, 2022, Appendix A.3). To simplify notation, we define $c_{k}(v_{lk})\coloneqq\|v_{lk}-\bar{d}_{k}\|^{p}-\epsilon^{p}$. We also note that $u_{lk}=Av_{lk}+b$. Then, we have
$\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})\\ \mbox{~{}~{}~{}~{}subject to}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\leq 0\end{array}$
$=\begin{array}[t]{ll}\underset{u_{11},\dots,u_{LK}\in S,\alpha\in\Gamma}{\sup}&\underset{\lambda\geq 0}{\inf}~{}\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk})\\ \end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}g_{l}(x,u_{lk})-\lambda\sum_{k=1}^{K}\sum_{l=1}^{L}\alpha_{lk}c_{k}(v_{lk}).\\ \end{array}$
We applied the Lagrangian in the first equality. Then, as the summation is over upper-semicontinuous functions $g_{l}(x,u_{lk})$ concave in $u_{lk}$, we applied the Von Neumann-Fan minimax theorem to interchange the inf and the sup. Next, we rewrite the formulation using an epigraph trick, and make a change of variables.
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{u_{11},\dots,u_{LK}\in S}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(x,u_{lk})-\lambda c_{k}(v_{lk}))\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{\alpha_{11}u_{11},\dots,\alpha_{LK}u_{LK}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}\alpha_{lk}(g_{l}(x,(\alpha_{lk}u_{lk})/\alpha_{lk})\\ &-\lambda c_{k}((\alpha_{lk}v_{lk})/\alpha_{lk}))\leq s_{k},\quad k=1,\dots,K.\end{array}$
In the last step, we rewrote $u_{lk}=(\alpha_{lk}u_{lk})/\alpha_{lk}$, $v_{lk}=(\alpha_{lk}v_{lk})/\alpha_{lk}$, and maximized over $\alpha_{lk}u_{lk}\in S^{\prime}$, where $S^{\prime}$ is the support transformed by $\alpha_{lk}$. See Wang et al., (2022) for details. Next, we make substitutions $h_{ij}=\alpha_{lk}u_{lk}$, $z_{ij}=\alpha_{lk}v_{lk}$, and define functions $g_{l}^{\prime}(x,h_{lk})=\alpha_{lk}g_{l}(x,h_{lk}/\alpha_{lk})$, $c_{k}^{\prime}(z_{lk})=\alpha_{lk}c_{k}^{\prime}(z_{lk}/\alpha_{lk})$.
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}~{}\underset{h_{11},\dots,h_{lk}\in S^{\prime}}{\sup}&\sum_{l=1}^{L}g_{l}^{\prime}(x,h_{lk})-\lambda c_{k}^{\prime}(z_{lk})\leq s_{k},\quad k=1,\dots,K\end{array}$
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}[-g_{l}^{\prime}(x,A\cdot+b)+\chi_{S^{\prime}}(A\cdot+b)+\lambda c_{k}^{\prime}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
For the new functions defined, we applied the definition of conjugate functions. Now, using the conjugate form $f^{\star}(y)=\alpha g^{\star}(y)$ of a right-scalar-multiplied function $f(x)=\alpha g(x/\alpha)$, we rewrite the above as
$=\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0}{\inf}&\sum_{k=1}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k}]^{\star}(0)\leq s_{k},\quad k=1,\dots,K.\end{array}$
We note that while the support of $h_{lk}$ is $S^{\prime}$, the support of $h_{lk}/\alpha_{lk}$, which is the input of $g_{l}$, is still $S$. Now, again using properties of conjugate functions, we note that:
$\alpha_{lk}[(-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda c_{k})]^{\star}(0)$ $=\alpha_{lk}\underset{\omega_{lk},h_{lk},y_{lk}}{\inf}([-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})$
$+\mathcal{I}_{A^{T}(h_{lk}+\omega_{lk})=y_{lk}}+\sigma_{S}(\omega_{lk})+[\lambda c_{k}]^{\star}(-y_{lk}))$
and
${}^{\star}(-y_{lk})=-y_{lk}^{T}\bar{d}_{k}+\phi(q)\lambda{\|}y_{lk}/\lambda{\|}^{q}_{*}+\lambda\epsilon^{p}.$
Substituting these back, we arrive at
$\begin{array}[t]{ll}\underset{\alpha\in\Gamma}{\sup}~{}\underset{\lambda\geq 0,h_{lk},\omega_{lk},s_{k}}{\inf}&\sum_{k}^{K}s_{k}\\ \mbox{subject to}&\sum_{l=1}^{L}\alpha_{lk}([-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})\\ &+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}+\phi(q)\lambda{\|}A^{T}(h_{lk}+\omega_{lk})/\lambda{\|}^{q}_{*}+\lambda\epsilon^{p})\leq s_{k},\\ &\hskip 56.9055pt\quad k=1,\dots,K,\quad j=1,\dots,J.\end{array}$
Taking the supremum over $\alpha$, noting that $\sum_{l=1}^{L}\alpha_{lk}=w_{k}$ for all $k$, we arrive at
$\begin{array}[t]{ll}\underset{\lambda\geq 0,z_{lk},\omega_{lk},s_{k}}{\inf}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\\ \mbox{s.t.}&[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K,\end{array}$
for which the final reformulation of (1) is
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&\lambda\epsilon^{p}+\sum_{k}^{K}w_{k}s_{k}\leq 0\\ &[-g_{l}]^{\star}(x,h_{lk})-b^{T}(h_{lk}+\omega_{lk})+\sigma_{S}(\omega_{lk})-(A^{T}(h_{lk}+\omega_{lk}))^{T}\bar{d}_{k}\\ &\hfill+\phi(q)\lambda\left\|A^{T}(h_{lk}+\omega_{lk})/\lambda\right\|^{q}_{*}\leq s_{k},\\ &\hfill l=1,\dots,L,\quad k=1,\dots,K.\end{array}$